### v2

# Fairness-in-the-Workflow: How Machine Learning Practitioners at Big Tech Companies Approach Fairness in Recommender Systems

Abstract.
Recommender systems (RS), which are widely deployed across high-stakes domains, are susceptible to biases that can cause large-scale societal impacts. Researchers have proposed methods to measure and mitigate such biases—but translating academic theory into practice is inherently challenging. RS practitioners must balance the competing interests of diverse stakeholders, including providers and users, and operate in dynamic environments. Through a semi-structured interview study (N=11), we map the RS practitioner workflow within large technology companies, focusing on how technical teams consider fairness internally and in collaboration with other (legal, data, and fairness) teams. We identify key challenges to incorporating fairness into existing RS workflows: defining fairness in RS contexts, particularly when navigating multi-stakeholder and dynamic fairness considerations. We also identify key organization-wide challenges: making time for fairness work and facilitating cross-team communication. Finally, we offer actionable recommendations for the RS community, including HCI researchers and practitioners.
recommender systems, algorithmic fairness, fairness in practice
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXconference: ACM CHI Conference on Human Factors in Computing Systems; April 13–17, 2026; Barcelona, Spainccs: Information systems Recommender systemsccs: Social and professional topics User characteristicsccs: Human-centered computing Empirical studies in HCI
1. Introduction
Recommender systems (RS) are inherently social and collaborative entities (Stray et al., 2024) which connect providers (e.g., sellers, content creators) with users (e.g., buyers, content consumers). RS are widely deployed across high-stakes domains including e-commerce (Gretzel and Fesenmaier, 2006; Zalmout et al., 2021) and social media (Lada et al., 2021; Belli et al., 2020; Anandhan et al., 2018)—but they are susceptible to biases that can cause large-scale psychological, ideological, and societal impacts within those domains (Li et al., 2021b; Chen et al., 2023; Kay et al., 2015; Sweeney, 2013). For example, Amazon’s e-commerce RS shows preference to Amazon’s own products even when competing products have superior reviews, resulting in an effective monopoly (Tate, 2023), and X’s social media RS disproportionately amplifies right-leaning political content, resulting in potential electoral impacts (Ye et al., 2025; Huszár et al., 2022). ML, HCI, and policy researchers have proposed a variety of methods to measure and mitigate such biases in order to improve fairness in RS (Feng and Shah, 2022; Koenecke et al., 2023; Smith et al., 2024; Abdollahpouri et al., 2019). However, appropriately translating any academic theory into industry application is inherently challenging because it requires navigating real-world constraints and balancing conflicting stakeholder interests (Holstein et al., 2019; Deng et al., 2022; Madaio et al., 2020; Beutel et al., 2019b; Veale et al., 2018; Metcalf et al., 2019; Robertson et al., 2022; Lee and Singh, 2021; Quiñonero Candela et al., 2023; Rieke et al., 2022; Smith et al., 2022). Incorporating fairness considerations into RS development is particularly challenging because RS operate within a landscape where fairness is not a static concept but a dynamic one, with meanings that shift across different contexts (Smith et al., 2024, 2025; Ekstrand et al., 2022; Smith et al., 2023b).
The complexity of RS is amplified by its operation in a multi-sided market, involving diverse stakeholders—both providers and end-users, each with heterogeneous make-ups (Burke, 2017; Abdollahpouri and Burke, 2019). RS practitioners who seek to incorporate fairness into their workflows therefore face the unique challenge of balancing multiple competing sets of stakeholder interests. Furthermore, RS practitioners operate subject to multiple ethical and regulatory constraints, such as the need to ensure the dissemination of trustworthy information, and sometimes must make tradeoffs between prioritizing fairness and those other considerations. RS practitioners also do not operate in a static environment. Their roles require not only identifying existing sources of unfairness, but anticipating emerging ones and mitigating their long-term effects, which can be exacerbated by feedback loops (Burke, 2017; Li et al., 2021b). This task demands a high level of foresight and adaptability—practitioners must be not only reactive, but also proactive in their approach to improving fairness in RS.
The ubiquity of RS today underscores the critical role that RS practitioners play in shaping society’s online interactions. Technical RS teams serve as the central hub where innovation, ethical considerations, and practical applications intersect. Here, the core challenges of fairness in RS—such as conflicting stakeholder interests and uncertain long-term impacts—are actively addressed and navigated. Hence, building on existing research, we explore the unique nuances of how machine learning (ML) practitioners at ‘big tech’ companies approach fairness in RS. In particular, we map the RS practitioner workflow, with a particular emphasis on how technical teams attempt to incorporate fairness (a) into their existing internal practices and (b) in collaboration with legal, data, and fairness teams. This allows us to understand how technical RS practitioners approach fairness in RS—a focus that is critical in understanding how to bridge academic theory on fairness in RS with actual practices. To do this, we conduct semi-structured interviews with RS technical practitioners who are integral members of ML teams at big tech companies (N=11). We examine what they have done to integrate fairness into their workflows, as well as the technical and organizational challenges they have faced to doing so. Technically, practitioners struggle to define fairness in the context of RS, especially because they must account for complexities inherent to RS, which include considering fairness across large numbers of (conflicting) stakeholder groups and working in dynamic settings with deployed systems that are susceptible to feedback loops. Organizationally, practitioners struggle with a lack of time to conduct fairness work and with developing a ‘fairness lingua franca’ that allows for clear communication across multidisciplinary teams. Our findings reveal opportunities for practitioners, ML researchers, and HCI researchers to improve fairness in RS by improving collaboration practices across industry teams and between industry and academia. Our key contributions are:
  • A map of the RS workflow, emphasizing how technical teams approach fairness (a) internally and (b) through collaborations with legal, data, and fairness teams
  • An overview of the key technical and organizational challenges to incorporating fairness into the RS workflow
  • Actionable steps towards addressing the challenges we identify by improving collaboration, communication, and documentation practices so that RS practitioners are better equipped to take lessons not only from industry fairness teams but also from academic researchers
We proceed as follows. First, we provide an overview of prior work on fairness considerations for RS, as well as prior work on the challenges inherent to translating academic theory into practice (§2). We describe our semi-structured interview and data analysis approach (§3). Then, we provide an overview of the RS workflow (§4). We describe technical challenges (§5) and organizational challenges (§6) to implementing fairness in RS workflows. We close with calls to action for the RS, ML fairness, and HCI communities (§7) as well as a discussion of the limitations of our work and potential paths forward (§8).
2. Related Work
Here, we provide a brief overview of the academic literature on ML fairness, with a particular focus on fairness for RS. Then, we discuss prior work on implementing fairness in deployed systems, with a focus on the challenges inherent to translating academic principles into practice and on the implications of ML fairness for industries and organizations.
2.1. Bias and Fairness in Machine Learning
Fairness and bias are closely related but distinct. Bias refers to systematic errors or deviations in data or models producing skewed outcomes for certain groups (Hutchinson and Mitchell, 2019). Fairness, by contrast, concerns the ethical acceptability of those biased outcomes (Corbett-Davies et al., 2023). We focus on fairness in this paper. [^1] We note that during interviews, interview participants sometimes used the terms ‘bias’ to refer to what we generally call ‘unfairness,’ or ‘debiasing’ to refer to what we generally call ‘improving fairness,’ so we use those terms when quoting participants directly. Because formal fairness definitions can be mutually incompatible, organizations must choose which principles to prioritize (Hutchinson and Mitchell, 2019; Ekstrand et al., 2022). For example, in a video-sharing platform, while one approach to fairness might involve using straightforward viewer metrics for ranking videos, an alternative approach might incorporate randomness in search rankings to counteract the “rich-get-richer” effect (Chen et al., 2023) wherein top-ranked videos get disproportionately more future views. The choice of how to define fairness often intersects with legal obligations (e.g., anti-discrimination laws) and business goals (Hutchinson and Mitchell, 2019; Ali et al., 2023). Our study builds on the fair ML literature by examining how fairness is approached in practice in RS workflows.
2.2. Fairness for Recommender Systems
RS pose unique fairness challenges due to multi-sided stakeholder needs (Chen et al., 2023; Wang et al., 2024b; Abdollahpouri and Burke, 2019; Burke, 2017). RS practitioners must balance the (potentially competing) interests of providers (e.g., sellers) and end-users (e.g., buyers)  (Burke et al., 2018; Singh and Joachims, 2018; Sonboli et al., 2021; Leonhardt et al., 2018; Li et al., 2021a; Stray et al., 2024; Mehrotra et al., 2018). For instance, a user-centric fairness approach might target equally accurate recommendations across demographics (Sonboli et al., 2021), while a provider-focused approach might seek equitable exposure for underrepresented creators (Abdollahpouri et al., 2019)—and these two interests can easily be in direct conflict with one another (Hutchinson and Mitchell, 2019; Sonboli et al., 2021; Ekstrand et al., 2022; Smith et al., 2023b). A system that prioritizes promoting content from historically marginalized providers might inadvertently serve users content that is misaligned with their previously stated interests (Smith et al., 2023b). And as Smith et al. (2024) point out, such a system can even have potentially negative effects for providers by exposing them to harassment from users who are outside of their preferred target audience. Complicating things further, a single platform may contain multiple types of providers, such as social media content creators and paid advertisers, whose interests may conflict as well. Despite the fact that RS are a commonly used technology, much ML fairness research has focused on classification problems (Barocas and Selbst, 2016; Chouldechova, 2017; Kleinberg et al., 2017; Dwork et al., 2012; Feldman et al., 2015; Zafar et al., 2017a; Corbett-Davies et al., 2017, 2023; Friedler et al., 2019; Kilbertus et al., 2017; Hardt et al., 2016b; Pleiss et al., 2017; Kamiran and Calders, 2009; Buolamwini and Gebru, 2018; Obermeyer et al., 2019; Friedler et al., 2021; Lee et al., 2024; Grgic-Hlaca et al., 2016; Hardt et al., 2016a; Sharifi-Malvajerdi et al., 2019; Zafar et al., 2017b; Borgs et al., 2019; Tal, 2023)—meaning that many existing fairness objectives and metrics cannot be directly applied to RS. Large-scale RS also raise unique challenges in practice related to sparse user-content interaction data, because users may only interact with a tiny subset of overall content (Li et al., 2021a; Kearns et al., 2019; Beutel et al., 2019b, a, 2018; Lum et al., 2022). Finally, as RS use user-content interaction data generated from the recommendation process to inform their future results, they risk feedback loops that amplify disparities (Keller, 2021). Our study contributes perspectives on implementing ML fairness in practice that are specific to RS, taking these unique challenges into account.
2.3. Fairness in Practice
In practice, fairness efforts do not occur in isolation. Rather, they interact—and may come into tension—with organizational priorities, such as maximizing profits. Acquiring organizational support to conduct responsible AI initiatives often requires linking fairness to explicit business or product goals (Scheuerman, 2024; Ali et al., 2023; Garcia-Gathright et al., 2018; Rakova et al., 2021; Veale et al., 2018; Wang et al., 2024a). Legal teams can also play a role in bringing about responsible AI efforts, although they often focus on privacy or liability rather than proactive fairness (Veale and Binns, 2017). However, regulations often vary across legal jurisdictions, complicating fairness efforts for multinational platforms that use RS. For example, RS in the EU are subject to the Digital Services Act (DSA), which authorizes authorities to enforce yearly assessments of online platforms regarding fairness issues, illegal content, and disruption to human rights (Madiega, 2020). The DSA does not apply to RS in the US (Fiordalisi, 2025); however, individual states have created their own fairness regulations, such as California’s 2022 state law AB 587, [^2] https://openstates.org/ca/bills/20212022/AB587/ which aims to provide transparency about large platform recommendations to avoid disinformation and hate speech. When fairness standards are not mandated by law, legal teams may actually hinder fairness efforts due to concerns that they could add to organizational liability (Costanza-Chock et al., 2022; Wright et al., 2024). If institutional support is lacking, employees may choose to “individualize risk” (Ali et al., 2023) by taking personal responsibility for incorporating fairness into their workflows. Prior work has repeatedly shown that these efforts can be challenging, as practitioners struggle to reconcile incompatible fairness metrics (Madaio et al., 2022; Moharana et al., 2025; Smith et al., 2025), work with messy real-world data (Cramer et al., 2019; Holstein et al., 2019; Rieke et al., 2022; Guerdan et al., 2025; Passi and Barocas, 2019; Sambasivan et al., 2021), and adapt general-purpose academic tools and metrics to their specific real-world needs (Deng et al., 2022; Harvey et al., 2025, 2024; Lee and Singh, 2021; Balayn et al., 2023; Kaur et al., 2020; Richardson et al., 2021; Ojewale et al., 2025). Many of these challenges are exacerbated by organizational challenges, such as communicating and collaborating across different teams and within cross-functional teams (Deng et al., 2023b). Researchers including Madaio et al. (2020), Deng et al. (2023a), and Costanza-Chock et al. (2022) have repeatedly found that a particular pain point when attempting to improve algorithmic fairness in practice is stakeholder involvement: practitioners struggle to involve the perspectives of end users in fairness assessments and improvements.
With the notable exception of Smith et al. (Smith et al., 2023b, a, 2024, 2025), most prior work on implementing fairness in practice has focused broadly on general-purpose ML or specifically on fairness in classification models, NLP models, or other non-RS applications of ML. However, we expect the issues raised by prior work to be even more pressing in the RS space, where fewer academic toolkits are available to scaffold fairness efforts, and where the interests of multiple stakeholders (providers and users) may be in direct competition. In fact, Smith et al. have previously identified tensions between competing stakeholders through interview studies with individuals who interact with RS but do not work directly on technical RS teams, including providers (content creators and dating app users) (Smith et al., 2024) and ML fairness experts (Smith et al., 2025). Most similar to our work, Smith et al. have also conducted interview studies with RS practitioners themselves: employees of an anonymous “commercial consumer-provider recommender system” (Smith et al., 2023a) and of a microlending platform (Smith et al., 2023b). Both studies explored challenges related to selecting fairness metrics in the context of RS, again identifying competing stakeholder interests as a key challenge (Smith et al., 2023b, a). In contrast, we interview technical RS practitioners across multiple big tech companies, and look beyond fairness metrics to how fairness is incorporated into RS workflows more broadly. In summary, we build on the extensive body of prior work by focusing explicitly on the RS workflow, with a particular spotlight not only on how RS practitioners across multiple big tech companies face technical challenges in the face of RS-specific complexity, but also how they navigate organizational challenges stemming from cross-team interactions with legal, data, and fairness practitioners.
3. Methods
To understand how RS practitioners approach fairness in practice, we conducted 11 semi-structured interviews over a three-week period in 2022. We interviewed RS practitioners from a variety of large companies that deploy highly visible RS.
Recruitment and Participants
We recruited participants via social media, direct emails to individuals in our networks or who we had identified based on publicly available information as working in RS, and snowball sampling. Participation was voluntary, and all participants provided informed consent prior to the start of interviews. Following the interviews, participants were compensated with Amazon gift cards. All interviews were conducted via video chat. Participants were given the option of either having interviews recorded to create audio transcripts, or having researchers take detailed notes during interviews. We provided the option for note-taking instead of recording as some practitioners expressed concerns about accidentally disclosing confidential company information, although no such disclosures ultimately occurred. In total, we recruited 11 ML practitioners (averaging 5.36 years of experience) across seven big tech companies. Practitioners held titles such as ‘Machine Learning Engineer,’ ‘Machine Learning Researcher,’ or ‘Scientist,’ and were actively involved in building RS. Each participant worked for a company that serves over one billion customers annually. Further details on participants are available in Table 1. This study was supervised by a university Institutional Review Board (IRB).
 Participant ID Years of Experience Engineering Responsibilities Research Responsibilities
 P1 3
P2 4
P3 4
P4 5
P5 6
P6 5
P7 4
P8 5
P9 7
P10 3
P11 12
Table 1. Participant Background Information.:
Interviews
Interviews ranged between 45 minutes and one hour. We began by asking ice-breaking questions regarding the practitioner’s background, including their years of experience, employer, and the RS that they built or worked with. Practitioners were instructed to be as detailed as possible but to avoid sharing any confidential information such as revenue or upcoming products. Next, we asked questions intended to zero in on practitioner workflows and approaches to fairness. We prepared a set of initial opening questions to encourage practitioners to share their general experiences and provide targets for follow-ups. We then proceeded to a set of follow-up questions about fairness metrics, practitioner time allocation, and cross-team collaboration based on earlier responses. Finally, we invited practitioners to comment on and rate their professional experiences in RS in terms of emotional support, work intensity, and stress level. Our semi-structured interview guide is provided in Appendix Table 3.
Analysis
We sought to understand how fairness is approached in practice by individuals and teams who develop and deploy large-scale RS. To do so, we analyzed responses on two levels: an affinity diagramming exercise to understand RS practitioner workflows, and a thematic analysis to understand the challenges faced by RS practitioners who seek to improve fairness in the systems that they build, deploy, and maintain.
Affinity Diagramming. First, we conducted an affinity diagramming exercise to map practitioners’ workflows based on interview notes and transcripts, following Yang et al. (2016). We focused on identifying commonalities, and also flagged outlier cases in our notes for further examination. We identified key stages practitioners shared in building RS. We also identified the core set of collaborators and external teams with whom RS practitioners interacted in order to construct a flow model of communications (Wixon et al., 1990). Data were analyzed during collaborative meetings between authors, with interviewers summarizing their sessions and outlining notes and/or transcripts relevant to each session under inspection. We then collaboratively sketched and refined models describing both the general flow of information and data during model development as well as the specific fairness approaches undertaken by practitioners. Findings are described in §4.
Thematic Analysis. Second, we conducted a thematic analysis to identify core challenges faced by practitioners (Braun and Clarke, 2019, 2006). We first converted scripts and notes into items on a Miro web board for aggregation and analysis. Authors conducted two coding passes through all quotes and reflections, developing a large set of initial theme groups. In the second round, authors aggregated thematic groups, merging similar themes and recategorizing elements through negotiation. Lastly, through another open discussion, authors broke down themes that had a significantly larger number of notes into finer granularity and placed any remaining unlabeled items into categories. When finalizing results, authors summarized the high-level themes identified and pulled representative quotes from the grouped items. Findings about technical challenges faced by RS practitioners are described in §5 and findings about organizational challenges arising from RS practitioner interactions with other teams are described in §6.
Positionality
We are a team of researchers who work on academic and industry teams, and who collectively have expertise in machine learning, human-computer interaction, and algorithmic fairness. Our backgrounds have certainly informed our approach to this study. Our experience on industry teams informed our understanding of the practical constraints that practitioners may face, which influenced the questions we asked in our semi-structured interviews. At the same time, our experience does not extend to working as RS practitioners within technology companies. While this positions us to critically assess RS workflows and challenges as outsiders, we do by definition lack ‘insider’ knowledge about what it is like to work on RS in practice. Finally, we are also all end-users of RS systems. We believe that this dual vantage point as both researchers and users shaped our understanding of the potential real-world impacts of RS, which in turn informed our approach to thematic analysis.
4. The RS Practitioner Workflow
Figure 1. Recommender System Practitioner Workflow. We visualize RS practitioners’ interactions with the data team managing and monitoring production data access, the legal team auditing and approving specific data access requests, and the fairness (or responsible AI) team providing fairness expertise. Distinct colors—blue for technical (RS practitioners), yellow for legal, red for data, and green for fairness—represent team inputs. Dotted arrows indicate occasional, rather than consistent, participation by a team in certain stages. Offline phases are denoted as gray boxes, while online phases are in purple boxes. Green circles indicate phases in which fairness considerations are commonly incorporated.:
In this section, we describe the RS practitioner workflow. We focus on how RS practitioners integrate fairness work into their labor, both internally (within technical teams) and in collaboration with legal, data, and fairness teams. These findings emerge out of the workflow-centric affinity diagramming exercise described in §3.
4.1. An Overview of the RS Practitioner Workflow
The RS practitioner workflow is summarized in Figure 1. Practitioners described workflows that evolved from an offline phase (gray boxes in Figure 1) into an online phase (purple boxes in Figure 1) throughout development. First, practitioners coordinate with the legal and data teams to receive access to necessary data and begin the data processing phase. The data team negotiates technical details such as API designs for collecting data and data cleaning, while the legal team reviews data access requests. Practitioners then spend much of their time on model construction and conducting rounds of offline model validation. All agreed that they invested more than half of their work time into these model-centric activities. For example, P5 reported, “This is the core of our daily work. I spent more than 60% time on model work.” Overall, the offline stages of the RS practitioner workflow align with the stages of general ML pipelines, e.g., as described by Suresh and Guttag (2021) and Black et al. (2023), as well as with preferences towards “model work” over “data work” that have been previously documented by Sambasivan et al. (2021).
After offline development, practitioners ship their model and move to an online testing phase, where real-world user behavior data and newly generated user-content interaction data are collected as users engage with recommendations. A fairness team, which all practitioners reported exists in their organizations under various names (such as ‘Responsible AI’), provides feedback to RS practitioners at this stage (see §4.2). Simultaneously, practitioners monitor the model’s performance metrics during the online model service phase, and might revisit the data pre-processing phase to change their training data (e.g., performing feature transformations or asking for more features).
4.2. How the RS Practitioner Workflow Incorporates Fairness
Fairness assessments and improvements occurred at several points throughout the RS workflow, corresponding to shifts to new development stages (green circles in Figure 1). These transition points were convenient locations to incorporate fairness because they often included interactions between different teams. Within these transition points, the overall process for incorporating fairness considerations into the RS practitioner workflow takes shape. It consists of three major phases, as depicted in Figure 2: the prototyping phase, the internal phase, and the collaborative phase. The prototyping phase co-occurs with the data processing phase of the RS workflow. In the data processing phase, RS practitioners interacted with legal and data teams to obtain access to training data. Concurrently, practitioners started the prototyping phase by hypothesizing about the sources of unfairness that might exist in the training data. Practitioners based hypotheses on their prior experience with RS, and did not generally cite academic research as a source for their hypotheses. Likewise, practitioners reported reusing fairness metrics that they had previously developed to assess training data in this stage. Later, during the offline model validation phase of the RS workflow, practitioners moved to the internal phase of the fairness process, wherein they validated the metrics developed during the prototyping phase and manually searched for unexpected results. Practitioners expressed that there are several rounds of iteration at this stage to find as many instances of potential unfairness as possible prior to model deployment.
Figure 2. Recommender System Practitioner Fairness Process. Technical teams incorporate fairness into their workflows in collaboration with legal, data, and fairness teams. :
After the model is shipped for online testing, the fairness process enters the collaborative phase. This phase is typically the first to include the fairness team, as well as the first phase to include data from real-world users of the RS. In the collaborative phase, the fairness team offers feedback based on their internal guidelines. Most practitioners reported that they had discretion as to whether to integrate feedback from the fairness team. Because practitioners felt that they had limited time and resources available, they sometimes attempted to weigh the relative impact of potential fairness issues against the work required to resolve them. As the owners of an RS within their organizations, practitioners reported numerous ongoing job requirements, such as monitoring the model’s online performance. Practitioners primarily reported being concerned about factors such as uptime, and did not consider fairness to reach a similar level of disruptive emergency necessitating urgent correction. When RS practitioners were able to address fairness issues in deployed models, they often had to integrate user feedback and data to seek out ‘hidden’ types of unfairness (i.e., those that they had not anticipated in the prototyping stage). This process is highly iterative and requires much manual analysis. Practitioners mentioned that, in a typical project, they eventually compose over five new fairness metrics tailored to the specific system and context, even though they typically started with pre-existing metrics from prior projects.
5. Technical Challenges Arising Within RS Practitioner Teams When Incorporating Fairness
Scope Type of Challenge Challenge
Technical
(Within RS Teams)
Defining Fairness Defining fairness in specific RS contexts, especially those not subject to pre-existing rules or regulation (11/11) Adapting pre-existing fairness metrics for RS (7/11), and managing the internal documentation involved in those adaptations (8/11)
Considering Multiple Types of Fairness Across Many Groups Managing (conflicting) multi-stakeholder interests (6/11) Prioritizing between different (non-conflicting) types of fairness (8/11) Handling high numbers of user groups (5/11)
Dynamic Fairness Considerations Developing ad hoc metrics to detect unexpected sources of bias, especially when managing real-world user interaction and feedback loops (10/11)
Organizational (Cross-Team) Lack of Time Finding time to work on fairness within technical teams (9/11) and in cross-team collaborations (6/11)
Cross-Team Communication Communicating and collaborating effectively with external teams (7/11)
Table 2. Challenges Faced by RS Practitioners when Incorporating Fairness into their Workflows. Challenges occur within RS teams as well as across teams and at the organization level. The number of practitioners who reported experiencing each type of challenge is shown in parentheses.:
In this section, we describe the key technical challenges faced by RS practitioners who seek to incorporate fairness considerations into their workflows. These findings emerge out of the challenge-centric thematic analysis described in §3. Table 2 details the technical challenges most commonly reported by practitioners.
5.1. RS Practitioners Struggle to Define Fairness in the Context of Recommender Systems
Defining Fairness.
Practitioners attempted to assess and improve fairness within two general categories. The first category is fairness as mandated by existing rules or regulations. For example, an RS that discriminated against job candidates from particular races would be considered unfair under US anti-discrimination law. The second category, encompassing all contexts not covered by existing rules or regulations, is perceived as more difficult and nuanced. One common frustration was that practitioners only have a vague goal (for example, not exposing content creators to accounts that might harass them), but are not armed with specific tools in their organization to reach the goal—or even a standardized method with which to approach the problem.
Adapting Pre-Existing Fairness Metrics and Navigating Documentation Debt.
While approaches for assessing and improving fairness have been proposed by fair ML literature (Chen et al., 2023), practitioners found them hard to directly adapt and ill-suited for their existing workflows. For example, P9 reported, “Even though we have a list of categories [over which to assess fairness] as you mentioned in research papers, we need to come up with a clear formula and implement it ourselves. It is really domain-specific.” Practitioners also struggled to build and re-use institutional knowledge about fairness in RS. This was due to the fact that RS practitioners faced documentation debt (Bender et al., 2021): prior fairness-related work at their organizations was not typically well-documented for future use. Half of practitioners interviewed therefore faced difficulties in re-using internal metrics that had been previously developed for fairness purposes. P7 shared, “For sure we can find their code, and comments, but without some context archived, for example, some documents, we are not sure whether we can use [them] directly.” Practitioners also felt ill-equipped to document their own fairness processes. As P3 stated, “Documentation is a good idea…however, I am not sure how specific of my debiasing should be included in the documentation.”
5.2. RS Practitioners Must Account for Multiple Types of Fairness Across Many Groups
Managing Conflicting Multi-Stakeholder Interests
Prior research emphasizes the dual role of recommender systems in simultaneously connecting and serving multiple stakeholders, including providers and end-users (Burke, 2017; Li et al., 2021b). This multifaceted interaction introduces complex challenges, especially in the context of fairness. The practitioners that we interviewed commonly made tradeoffs, often prioritizing one user group over others. However, deciding how to make these tradeoffs was a challenge, as highlighted by P9: “Tradeoffs of a different group of users, or accuracy of the recommendations are hard to be decided,” and P1, who pointed out a lack of a “compelling [Standard Operating Procedure] for those decisions.” Making tradeoffs, while common, may not always be necessary. Recent academic findings suggest that multi-objective optimization might not always adversely affect particular user groups (Ekstrand et al., 2022), indicating a disconnect between industry practices and academic research.
Prioritizing Between Different (Non-Conflicting) Types of Fairness.
Practitioners felt that there was more of a “what to prioritize” question in RS compared with traditional predictive models. In a traditional ML pipeline, fairness assessment focuses on one particular target (e.g., demographic parity). However, the practitioners that we interviewed expressed that they have to balance multiple fairness-related considerations. For example, one participant needed to consider demographic parity for job seekers in one region—and simultaneously, ensure that advertisers across multiple regions were adhering to different sets of fairness-related regulatory requirements. Importantly, these fairness considerations were not necessarily in conflict with one another. However, because practitioners had limited time and resources available, they nevertheless needed to choose certain fairness ideals to prioritize before turning to others. P3 pointed out: “Debiasing for multiple purposes is crucial in ensuring a fair and unbiased platform for all stakeholders. It’s not just about the debiasing of content creators based on region, gender, and religion…but also about addressing other factors that can impact the platform…For instance, we need to make sure that products…are supported fairly [e.g., that products that are artificially cheap because of government subsidies designed to gain market dominance are downranked on our RS platform].”
Handling High Numbers of User Groups
A significant challenge identified by practitioners involves ensuring fairness across a wide range of groups within RS platforms. Current metrics often fail to address the complexities arising from the large number of groups involved (and the ensuing intersectional groups). P2 described this challenge: “It’s hard to manage disparities between the numerous groups. With thousands of groups to consider, it becomes nearly impossible to ensure uniformity across all.” This sentiment was echoed by P9, who noted, “Most metrics are designed with the assumption of a few groups…They don’t cater to intra-group-level debiasing, making our adaptation efforts cumbersome.” Additionally, P7 highlighted the computational difficulties in handling sparse datasets resulting from subgroup division: “Most of the training data is too sparse, and it is hard for us to enforce the disparities between groups. I cannot look through thousands of groups to make sure that holds.”
5.3. Fairness Considerations are Complicated by Dynamic RS Environments
Developing Ad Hoc Metrics to Detect Unexpected Instances of Unfairness.
Practitioners engaged in building RS often create bespoke metrics in collaboration with their teams. P2 and P3 highlighted the personalized nature of these metrics, with P2 noting, “We proposed our own definitions of bias,” and P3 adding, “Most of the metrics we developed are intuition-based.” Such approaches underscore the reliance on practitioner expertise and intuition in assessing and improving fairness, as echoed by P6: “In general, I would like to write some hand-crafted codes to validate my hypothesis. Experience is a key to debiasing sometimes,” despite the availability of internal fairness tools. Nevertheless, practitioners acknowledged that their initial metrics often fell short in detecting unfairness, and they therefore engaged in iterative refinement of those metrics based on user feedback:
[P1]: It is not an easy job…so we need to start from somewhere. After doing some debiasing based on your experience, we look forward to getting feedback from real-world users to further debiasing.
[P7]: It sounds a bit contradictory. We have a model first and keep debiasing the model based on our user’s new data. It although risks bringing in some confirmatory bias, it works for now.
Managing Real-World User Interactions and Feedback Loops
Practitioners also shed light on the challenges of assessing and improving fairness in online (deployed), as compared to offline, RS. This challenge is particularly acute in RS given the volume of data, the subtleties of human interactions, and the potential for user feedback loops to amplift latent sources of unfairness. P1 and P2 emphasized the difficulty in uncovering unexpected fairness issues in deployed systems, with P1 stating, “As the bias definitions are so different, and there are so many different user groups, those biases are hiding well behind the recommender systems that we built,” and P2 reflecting, “Most of the time, we have to look carefully to find those hidden biases. It is not fun.” Despite finding academic metrics less directly applicable (see §5.1), practitioners acknowledged the value of guidelines from research in providing a foundation and confidence in their decision-making processes. For example, P7 stated, “Though sometimes [checklists] are not directly related, I can still check those and they help me think through my reasoning. It made me somehow more confident in my decision.”
6. Organizational Challenges Arising Across Teams When Incorporating Fairness
In this section, we describe the key organizational challenges faced by RS practitioners who seek to incorporate fairness considerations into their workflows. As with the technical challenges outlined in §5, these findings emerge out of the challenge-centric thematic analysis described in §3. Table 2 details the organizational challenges most commonly reported by practitioners.
6.1. RS Practitioners Lack Time to Engage in Collaborative Fairness Work
Practitioners universally described their workflows as highly iterative and time-sensitive. As a result, most felt they had a very limited amount of time to focus on fairness, either internally or in collaborations with the fairness team. Nine reported that they spent roughly 5% percent of their work time on fairness-related tasks, and two said they spent up to 10% of their time. As P7 put it, “The negative effect spurred from the platform I worked for before is not so ‘bad.’ We don’t have the incentive to invest half of our time…to comprehensively debias.” When prompted to reflect on this allocation, most practitioners indicated that, while allocating more time for fairness-related tasks could be helpful, they were neutral about doing so. P1 reported, “Even though I wish I could have more time on this, I understand why this proportional[ly] makes sense. From the perspective of our org, prioritizing efforts on improving ‘important’ metrics brings more impact.” In other words, fairness issues did not rise to the level of a disruptive emergency, and thus were not prioritzed by management in evaluating project outcomes.
6.2. RS Practitioners Lack a ‘Fairness Lingua Franca’ to Facilitate Cross-Team Communication
Engaging with external fairness teams revealed a nuanced challenge: attaining mutual understanding and converging perspectives. Practitioners valued the fairness teams’ feedback but sometimes struggled with its practical integration due to a few key complexities. Firstly, the inherent structure of RS as complex deep learning models sometimes made feedback from fairness teams difficult to implement. P9 observed, “sometimes…[a] challenge is fundamentally unsolvable, rooted in the deep learning models.” P1 shared, “Their feedback could be something I agree with, but I cannot solve.” Secondly, the multidisciplinary nature of fairness teams led them to use terminology that was not readily understood by technical teams, affecting inter-team communication. P9 commented on the influence of varied backgrounds and perspectives on understanding: “sometimes I don’t understand some phrases they used in their feedback, for example, ‘inductive bias from recommender systems,’ and need to ask for more explanations.” P6 added, “Some of my colleagues are from the field of psychology and communication, and some of those terms I didn’t know before.” Thirdly, practitioners desired clearer guidance from fairness teams, particularly regarding model development. P7 said, “Once they mentioned a certain aspect of my model might introduce bias, I wished to understand the reasoning.” Finally, fairness teams were often involved only after models were deployed for online testing. P9 suggested, “Early involvement of the fairness team would have facilitated a closer working relationship.” P2 noted, “Sometimes there’s a lag in syncing up; feedback might refer to previous versions of our recommender system designs.”
These challenges starkly contrast with how RS practitioners interact with legal teams. Due to the sensitive nature of end-user data, practitioners regularly interact with their legal team to confirm the scope of data that can be gathered. With increasing surveillance from authorities and global concerns over privacy on RS, one might expect the relationship between practitioners and internal policymakers to be fraught. However, practitioners reported that communication with legal teams was clear and direct. Per P2, “If they said ‘no,’ then it is a clear ‘no,’ and we know the reasons and risks behind their decision.” Similarly, P3 stated, “Their guidelines are very clear, particularly when compared with fairness guidelines, which we don’t really have.” Clear and open guidelines from legal teams made it easier for practitioners to work with them, and most practitioners acknowledged the role of the legal team in avoiding risks. For instance, P2 mentioned that the “legal team serves as the gatekeeper to keep us away from legal risks, and help us comply well with our company’s vision and values.” P1 summarized the issue succinctly: “Our collaborations [with fairness teams] could be streamlined if there were guidelines as clear as those from legal teams.”
7. Calls to Action
In this section, we map the pain points uncovered in our interviews to calls for action, targeted towards RS and fairness communities alike.
7.1. Building Institutional Knowledge Through Documentation and Collaboration
A key challenge reported by RS practitioners was adapting fairness metrics that had been previously developed at their organizations for new fairness work. Teams are composed of individuals from a wide variety of personal backgrounds, numerous project stages are spread over time, and systems are improved dynamically after deployment, when fairness effects may be hard to predict and can potentially self-reinforce in feedback loops. Due to these challenges, documenting prior fairness efforts is critical for building institutional fairness knowledge. However, as multiple practitioners pointed out, while there are standard practices for documenting and approving code changes, there is comparably little infrastructural support for documenting approaches to assess and improve fairness. Researchers have previously called for such fairness documentation (Heger et al., 2022), e.g., in the style of datasheets for datasets (Gebru et al., 2021; Papakyriakopoulos et al., 2023). An internal template for fairness documentation for RS can help to improve the process in the immediate term and also build institutional knowledge.
Considering that RS involve massive amounts of content, large user bases, rapidly changing environments, and large teams of developers, it is surprising that the practitioners we interviewed did not identify stronger support for maintaining team cohesion and situational awareness during project development. While collaboration was explicitly cited as a stage of the fairness process by practitioners, practitioners did not report receiving specific technical or procedural support to facilitate cross-team collaboration. We believe there is a need for research into best practices for collaboration and information exchange during RS development to better support cross-team fairness efforts.
7.2. Encouraging Researcher and Organizational Investment in Multi-Stakeholder Fairness
To tackle the challenges involved in improving fairness in RS, organizations must invest intellectual resources into researching, or working to implement academic research on, several key areas. Firstly, it is necessary to develop more comprehensive optimization approaches, considering multiple objectives, to balance (or avoid) tradeoffs between different stakeholder groups. Secondly, further investigations into group-wise fairness performance comparisons should address the “too many groups” challenge in RS by making it easier to assess fairness across a very high numbers of groups. Similarly, there should also be an increased focus on intra-group fairness to align with multi-faceted and intersectional conceptions of diversity in the real world. Additionally, researchers and organizations should work to develop standardized guidelines for making tradeoff decisions to scaffold decision-making when fairness tradeoffs between stakeholder groups is unavoidable. This would help avoid the ad hoc and potentially arbitrary decision-making that currently occurs when RS practitioners are forced to make these decisions in real time as tradeoffs arise. Guidelines should be developed in collaboration with diverse stakeholder groups, and should encourage practitioners to question, rather than reinforce, existing power dynamics on RS platforms.
7.3. Integrating Fairness Considerations into Workflows Early and Often
On an organizational level, we found that practitioners encountered competing incentives for integrating fairness into their workflows. While practitioners’ primary job responsibilities were to create and manage RS, every practitioner interviewed felt both a personal interest and an explicit requirement to assess and improve fairness in their models. At the same time, however, we noticed many cases where practitioners had to backtrack or iterate on deployed models in order to mitigate unexpected instances of unfairness. Had they invested additional time during project formation and received support from fairness teams earlier, perhaps such issues could have been detected when changes would have been easier to make. Our workflow models (Figures 1 and 2) highlight the opportunity for early engagement with fairness teams, at transition points in the workflow and during phases of the RS workflow that align with phases of the fairness process. We also call for additional research into the organizational context of RS in order to identify organizational barriers beyond the scope of our investigation.
7.4. Developing a ‘Fairness Lingua Franca’
Across Industry Teams
We identified cases where poor communication between fairness teams and technical practitioners led to confusion and even deprioritization of fairness efforts. While legal teams offered explicit guidance and directives to practitioners, the nature of fairness feedback—reliant on nuance and not necessarily immediately applicable to complex models—proved much harder to parse and operationalize in systems. Combined with the fast-paced nature of RS development, this led to disconnects between technical practitioners and fairness teams despite mutual agreement on the importance of fairness. We believe that these disconnects could have been anticipated. Practitioners reported that fairness teams came from multiple disciplines, including some (like psychology and communication) which may not share common terminology with ML. RS practitioners lacked the broad interdisciplinary expertise represented on fairness teams, resulting in a disconnect in fully understanding the implications and context of fairness feedback. The dynamic environment of RS exacerbated this gap, leading practitioners to optimize by solving easy issues while neglecting other—potentially more serious—issues. We found that the academic literature was often viewed as a reliable, albeit not universally applicable, resource by both practitioners and fairness teams. We therefore posit that HCI and ML researchers might also help with the necessary ‘translation’ work to provide useful vocabulary and common ground for practitioners. Likewise, embedding fairness practices more deeply in technical teams through additional roles, building domain expertise in mutual fields, and pursuing techniques for more consistent documentation and information flow (especially at the beginning of the RS workflow) might also help to establish common ground.
Across Industry and Academia
We found that knowledge from research communities not only shaped many practitioners’ awareness of fairness but also gave them some basic principles to follow. Yet practitioners also regularly developed their own ad hoc fairness metrics instead of using those described in academic literature. This creates a patchwork of practices that at times adheres to academic recommendations and at other times drifts far from them. While practitioners do not necessarily seek to disregard academic best practices, practical constraints sometimes make it hard to follow, or even known about, them. We call for greater collaboration between industry and academia, acknowledging that accomplishing this goal is challenging and limited by many factors beyond the scope of our study. However, there are clear benefits to both sides: industry practitioners may benefit from investigating novel approaches, while research communities can potentially gain exposure to new use cases and novel empirical evidence.
8. Discussion
8.1. The Path Forward: RS Practitioners Emphasize Benefits of Fairness Work Despite Challenges
Despite the aforementioned challenges, we have reason to be optimistic: RS practitioners consistently ascribe importance to fairness work. In our semi-structured interviews, we asked practitioners to describe their emotional stress and satisfaction level in their daily work. While most practitioners reported that balancing daily work and fairness considerations was somewhat stressful, most also believed that fairness is an important and necessary part of their work. P9 described incorporating fairness as “…tiring but really rewarding.” Practitioners pointed out that fairness felt particularly important given the uniquely large impact that RS play on society. P1 notes: “It is tiring; however, I believe this is the right thing to do. It is important to remove bias as much as possible as our system serves millions of users worldwide. Discriminating behaviors imposed by our system might greatly affect their life.” Practitioners expressed a desire to improve the world through their systems and identified strong intrinsic motivations for improving fairness, even if it is arduous. Some practitioners cited public and academic discourse on fairness as motivating; for instance, P1 mentioned, “Thorough discussions from existing papers, you know, like COMPAS, about how unfairness affects other people’s life greatly shaped my view on this. As I have the chance to repair this, I will do my best—though it is not easy.”
8.2. Limitations and Ethical Considerations
Limitations
Our study is subject to a number of limitations. First and foremost, our findings and discussions are based on a limited sample of practitioners from seven highly visible companies. Although our sample size is typical of HCI interview studies (Caine, 2016), and although we intentionally targeted practitioners working on very large-scale RS with high societal impact, this limits the kind of evidence that we were able to gather. In particular, our findings may not generalize to RS practitioners working on smaller-scale RS. Similarly, this paper focuses on machine learning practitioners working on RS who are leading engineering or research efforts; we did not interview members of legal, data, or fairness teams. The desires we uncovered may also not be fully actionable: there is a fundamental tension between the desire for fairness rules that function similarly to rigid and straightforward legal and data governance requirements, and the ability actually to implement such rules in the wild (Ali et al., 2023). Finally, our semi-structured interviews might be limited by the remote setting, potential risks of exposing corporate secrets, or participant self-presentation maintenance. While we found participants to generally be quite frank about their experiences, it is possible that they provided a more critical or a rosier picture of the environment due to their perceived expectations.
Ethical Considerations
Our primary ethical consideration in conducting this study was related to the privacy and confidentiality of our interview subjects. In order to limit the possibility of re-identification, we provide only limited information about interview participants and their employers. We strictly adhered to our IRB guidelines during the interview and data analysis process by asking participants not to share sensitive information, de-identifying interview transcripts, and securely storing interview data.
Adverse Impacts
While this study aims to foster understanding and improve fairness in RS, we acknowledge the potential for adverse impacts once published. The findings could unintentionally influence RS practitioners to adopt fairness practices that may not be universally applicable, leading to oversimplification of complex issues. Furthermore, there is a risk that the emphasis on certain strategies might overshadow other equally important aspects of fairness in RS. We recognize the importance of ongoing critical evaluation and discourse in the community to mitigate these potential negative impacts and encourage a diverse range of perspectives and solutions in addressing structural issues of fairness in recommender systems.
9. Conclusion
In this work, we investigated fairness practices in RS through a semi-structured interview study of 11 technical practitioners who are building recommender systems. We constructed a process-level map of RS practitioners’ daily workflow, focusing on how they interact with legal, data, and fairness teams at their organizations and on how they incorporate fairness into their existing practices. We identified key technical challenges to incorporating fairness, such as defining fairness in the context of RS and navigating the complexity introduced by RS; namely, multi-stakeholder fairness considerations and dynamic environments. We also identified organizational challenges resulting from cross-team interactions, including a lack of time to conduct fairness work and an communication challenges across teams stemming from a lack of a ‘fairness lingua franca.’ We suggest that building institutional knowledge by improving documentation and collaboration practices, encouraging researcher and organizational investment into multi-stakeholder fairness approaches, integrating fairness considerations into RS workflows early and often, and developing a ‘fairness lingua franca’ can lead to improvements in RS fairness.
Acknowledgements.
This work was supported by NSF grant IIS-1850195. We would like to thank the associate chairs and anonymous reviewers for their invaluable feedback. We also would like to thank Dataprep team from Simon Fraser University for the valuable input of mapping of common functions and user intentions. All opinions, findings, and conclusions in this paper are those of the authors and do not necessarily reflect the views of the funding agencies.

### v3

@@
-Fairness-in-the-Workflow: How Machine Learning Practitioners at Big Tech Companies Approach Fairness in Recommender Systems
+arXiv:2505.19441v3[cs.HC] 27 Feb 2026
 Abstract.
@@
 Researchers have proposed methods to measure and mitigate such biases—but translating academic theory into practice is inherently challenging.
-RS practitioners must balance the competing interests of diverse stakeholders, including providers and users, and operate in dynamic environments.
-Through a semi-structured interview study (N=11), we map the RS practitioner workflow within large technology companies, focusing on how technical teams consider fairness internally and in collaboration with other (legal, data, and fairness) teams.
-We identify key challenges to incorporating fairness into existing RS workflows: defining fairness in RS contexts, particularly when navigating multi-stakeholder and dynamic fairness considerations.
+Through a semi-structured interview study (N=11), we map the RS practitioner workflow within large technology companies, focusing on how technical teams consider fairness internally and in collaboration with legal, data, and fairness teams.
+We identify key challenges to incorporating fairness into existing RS workflows: defining fairness in RS contexts, balancing multi-stakeholder interests, and navigating dynamic environments.
 We also identify key organization-wide challenges: making time for fairness work and facilitating cross-team communication.
-Finally, we offer actionable recommendations for the RS community, including HCI researchers and practitioners.
+Finally, we offer actionable recommendations for the RS community, including practitioners and HCI researchers.
 recommender systems, algorithmic fairness, fairness in practice
-copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXconference: ACM CHI Conference on Human Factors in Computing Systems; April 13–17, 2026; Barcelona, Spainccs: Information systems Recommender systemsccs: Social and professional topics User characteristicsccs: Human-centered computing Empirical studies in HCI
+journalyear: 2026copyright: ccconference: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems; April 13–17, 2026; Barcelona, Spainbooktitle: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems (CHI ’26), April 13–17, 2026, Barcelona, Spaindoi: 10.1145/3772318.3791347isbn: 979-8-4007-2278-3/2026/04ccs: Information systems Recommender systemsccs: Social and professional topics User characteristicsccs: Human-centered computing Empirical studies in HCI
 1. Introduction
-Recommender systems (RS) are inherently social and collaborative entities (Stray et al., 2024) which connect providers (e.g., sellers, content creators) with users (e.g., buyers, content consumers).
-RS are widely deployed across high-stakes domains including e-commerce (Gretzel and Fesenmaier, 2006; Zalmout et al., 2021) and social media (Lada et al., 2021; Belli et al., 2020; Anandhan et al., 2018)—but they are susceptible to biases that can cause large-scale psychological, ideological, and societal impacts within those domains (Li et al., 2021b; Chen et al., 2023; Kay et al., 2015; Sweeney, 2013).
+Recommender systems (RS) are inherently social and collaborative entities (Stray et al., 2024; Li et al., 2025) which connect providers (e.g., sellers, content creators) with users (e.g., buyers, content consumers).
+RS are widely deployed across high-stakes domains including e-commerce (Gretzel and Fesenmaier, 2006; Zalmout et al., 2021) and social media (Lada et al., 2021; Belli et al., 2020; Anandhan et al., 2018)—but they are susceptible to biases that can cause large-scale psychological, ideological, and societal impacts (Li et al., 2021b; Chen et al., 2023; Kay et al., 2015; Sweeney, 2013).
 For example, Amazon’s e-commerce RS shows preference to Amazon’s own products even when competing products have superior reviews, resulting in an effective monopoly (Tate, 2023), and X’s social media RS disproportionately amplifies right-leaning political content, resulting in potential electoral impacts (Ye et al., 2025; Huszár et al., 2022).
@@
 RS practitioners also do not operate in a static environment.
-Their roles require not only identifying existing sources of unfairness, but anticipating emerging ones and mitigating their long-term effects, which can be exacerbated by feedback loops (Burke, 2017; Li et al., 2021b).
-This task demands a high level of foresight and adaptability—practitioners must be not only reactive, but also proactive in their approach to improving fairness in RS.
+Their roles require not only identifying existing sources of unfairness, but anticipating emerging ones and mitigating their long-term effects, which can be exacerbated by user feedback loops (Burke, 2017; Li et al., 2021b, 2025).
+This task demands a high level of foresight and adaptability—practitioners must be proactive in their approach to improving fairness in RS.
 The ubiquity of RS today underscores the critical role that RS practitioners play in shaping society’s online interactions.
 Technical RS teams serve as the central hub where innovation, ethical considerations, and practical applications intersect.
-Here, the core challenges of fairness in RS—such as conflicting stakeholder interests and uncertain long-term impacts—are actively addressed and navigated.
+Here, the core challenges of fairness in RS are actively addressed and navigated.
 Hence, building on existing research, we explore the unique nuances of how machine learning (ML) practitioners at ‘big tech’ companies approach fairness in RS.
-In particular, we map the RS practitioner workflow, with a particular emphasis on how technical teams attempt to incorporate fairness (a) into their existing internal practices and (b) in collaboration with legal, data, and fairness teams.
-This allows us to understand how technical RS practitioners approach fairness in RS—a focus that is critical in understanding how to bridge academic theory on fairness in RS with actual practices.
-To do this, we conduct semi-structured interviews with RS technical practitioners who are integral members of ML teams at big tech companies (N=11).
+In particular, we map the RS practitioner workflow, with an emphasis on how technical teams attempt to incorporate fairness (a) into their existing internal practices and (b) in collaboration with legal, data, and fairness teams.
+This allows us to understand how technical practitioners approach fairness in RS—a focus that is critical in understanding how to bridge academic theory with actual practices.
+To do this, we conduct semi-structured interviews with technical RS practitioners who are integral members of ML teams at big tech companies (N=11).
 We examine what they have done to integrate fairness into their workflows, as well as the technical and organizational challenges they have faced to doing so.
@@
 Organizationally, practitioners struggle with a lack of time to conduct fairness work and with developing a ‘fairness lingua franca’ that allows for clear communication across multidisciplinary teams.
-Our findings reveal opportunities for practitioners, ML researchers, and HCI researchers to improve fairness in RS by improving collaboration practices across industry teams and between industry and academia.
+Our findings reveal opportunities for practitioners and HCI researchers to improve fairness in RS.
 Our key contributions are:
@@
   • An overview of the key technical and organizational challenges to incorporating fairness into the RS workflow
-  • Actionable steps towards addressing the challenges we identify by improving collaboration, communication, and documentation practices so that RS practitioners are better equipped to take lessons not only from industry fairness teams but also from academic researchers
+  • A set of actionable recommendations to better integrate fairness into RS workflows, highlighting both where big tech companies and practitioners are responsible for implementing evidence-based changes as well as where novel HCI research can support real-world fairness efforts
 We proceed as follows.
-First, we provide an overview of prior work on fairness considerations for RS, as well as prior work on the challenges inherent to translating academic theory into practice (§2).
-We describe our semi-structured interview and data analysis approach (§3).
+First, we provide an overview of prior work on fairness considerations for RS and on the challenges inherent to translating academic theory into practice (§2).
+We describe our semi-structured interview and analysis approach (§3).
 Then, we provide an overview of the RS workflow (§4).
 We describe technical challenges (§5) and organizational challenges (§6) to implementing fairness in RS workflows.
-We close with calls to action for the RS, ML fairness, and HCI communities (§7) as well as a discussion of the limitations of our work and potential paths forward (§8).
+We make calls to action for RS practitioners and HCI researchers (§7) and close with a discussion of the limitations of our work and potential paths forward (§8).
 2. Related Work
-Here, we provide a brief overview of the academic literature on ML fairness, with a particular focus on fairness for RS.
-Then, we discuss prior work on implementing fairness in deployed systems, with a focus on the challenges inherent to translating academic principles into practice and on the implications of ML fairness for industries and organizations.
-2.1. Bias and Fairness in Machine Learning
+Here, we provide a brief overview of the academic literature on ML fairness, with a focus on fairness for RS.
+We also discuss prior work on implementing fairness in deployed systems, with a focus on the challenges inherent to translating academic principles into practice.
+2.1. Fairness in Machine Learning
 Fairness and bias are closely related but distinct.
-Bias refers to systematic errors or deviations in data or models producing skewed outcomes for certain groups (Hutchinson and Mitchell, 2019).
+Bias refers to systematic errors or deviations in data or models that produce skewed outcomes for certain groups (Hutchinson and Mitchell, 2019).
 Fairness, by contrast, concerns the ethical acceptability of those biased outcomes (Corbett-Davies et al., 2023).
 We focus on fairness in this paper.
-[^1] We note that during interviews, interview participants sometimes used the terms ‘bias’ to refer to what we generally call ‘unfairness,’ or ‘debiasing’ to refer to what we generally call ‘improving fairness,’ so we use those terms when quoting participants directly.
-Because formal fairness definitions can be mutually incompatible, organizations must choose which principles to prioritize (Hutchinson and Mitchell, 2019; Ekstrand et al., 2022).
-For example, in a video-sharing platform, while one approach to fairness might involve using straightforward viewer metrics for ranking videos, an alternative approach might incorporate randomness in search rankings to counteract the “rich-get-richer” effect (Chen et al., 2023) wherein top-ranked videos get disproportionately more future views.
-The choice of how to define fairness often intersects with legal obligations (e.g., anti-discrimination laws) and business goals (Hutchinson and Mitchell, 2019; Ali et al., 2023).
-Our study builds on the fair ML literature by examining how fairness is approached in practice in RS workflows.
-2.2. Fairness for Recommender Systems
-RS pose unique fairness challenges due to multi-sided stakeholder needs (Chen et al., 2023; Wang et al., 2024b; Abdollahpouri and Burke, 2019; Burke, 2017).
-RS practitioners must balance the (potentially competing) interests of providers (e.g., sellers) and end-users (e.g., buyers)  (Burke et al., 2018; Singh and Joachims, 2018; Sonboli et al., 2021; Leonhardt et al., 2018; Li et al., 2021a; Stray et al., 2024; Mehrotra et al., 2018).
+[^1] We note that during interviews, participants sometimes used the terms ‘bias’ to refer to what we generally call ‘unfairness,’ or ‘debiasing’ to refer to what we generally call ‘improving fairness,’ so we use those terms when quoting participants directly.
+Fairness is an essentially contested concept, meaning that there is no single agreed-upon approach for defining fairness (Jacobs and Wallach, 2021).
+A dominant fairness paradigm focuses on the outcomes of technical systems: systems are group fair to the extent that they produce similar outcomes for different groups of individuals, and individually fair to the extent that they treat ‘similar’ individuals similarly (Dwork et al., 2012; Barocas and Selbst, 2016).
+Although these approaches sound simple, they are anything but: group fairness can be defined using multiple metrics, many of which are impossible to simultaneously satisfy, and individual fairness requires a way of determining how similar any two individuals are, which is often infeasible in practice (Chouldechova, 2017; Kleinberg et al., 2017; Zemel et al., 2013).
+An alternative paradigm draws on the theory of procedural justice (Rawls, 1999), arguing that systems are fair to the extent that they follow fair processes—although what makes a process fair is not settled, and may involve anything from only using certain features (Grgic-Hlaca et al., 2016) to incorporating multiple stakeholder groups in participatory system development (Lee et al., 2019).
+Finally, fairness is but one component of the broader field of ethical or responsible AI.
+Thus, the choice of how to define fairness often intersects with other considerations.
+These may include legal obligations (e.g., anti-discrimination laws) (Hildebrandt, 2022; Magrani and da Silva, 2024), explainability (Kunkel et al., 2019), user privacy considerations (Sun et al., 2023), and the extent to which systems facilitate appeal and recourse of algorithmic decisions (Jin and Salehi, 2024).
+Ultimately, organizations must choose which principles to prioritize, and the choice is far from simple (Hutchinson and Mitchell, 2019; Ekstrand et al., 2022).
+Fairness for Recommender Systems
+Defining fairness is particularly challenging in the context of RS because much ML fairness research has focused on classification problems, meaning that many existing fairness objectives and metrics may not be directly applicable to RS.
+For example, although fairness in classification problems often focuses on outcomes, what an outcome means in the context of RS is not necessarily clear: outcomes can be thought of in terms of not only allocations (e.g., whose content gets recommended) but also representations (e.g., whether recommended content contributes to the stereotyping, erasure, or demeaning of social groups) (Blodgett et al., 2020).
+Nevertheless, researchers have proposed a variety of approaches for measuring and improving fairness in recommendations and ranking (Zehlike et al., 2017; Asudeh et al., 2019; Biega et al., 2018; Ali et al., 2019; Beutel et al., 2019a; Diaz et al., 2020; Ekstrand et al., 2018; Geyik et al., 2019; Li et al., 2021a; Zehlike and Castillo, 2020; Singh and Joachims, 2018; Steck, 2018; Yang and Stoyanovich, 2017; Rastogi and Joachims, 2024).
+[^2] For comprehensive overviews of fairness in RS, we direct the reader to Ekstrand et al.
+(2022), Patro et al.
+(2022), and Zehlike et al.
+(2022a, b).
+For a survey of fair ranking metrics, we point to Smith et al.
+(2023a).
+As discussed above, defining fairness (e.g., by choosing a fairness metric) is challenging in ML generally.
+It is uniquely challenging in the context of RS because RS require balancing the (potentially competing) interests of multiple stakeholders who are dynamically interacting with each other and the RS platform at massive scales (Chen et al., 2023; Wang et al., 2024b; Abdollahpouri and Burke, 2019; Burke, 2017).
+In particular, RS practitioners must consider the interests of providers (e.g., sellers) and end-users (e.g., buyers)  (Burke et al., 2018; Singh and Joachims, 2018; Sonboli et al., 2021; Leonhardt et al., 2018; Li et al., 2021a; Stray et al., 2024; Mehrotra et al., 2018).
 For instance, a user-centric fairness approach might target equally accurate recommendations across demographics (Sonboli et al., 2021), while a provider-focused approach might seek equitable exposure for underrepresented creators (Abdollahpouri et al., 2019)—and these two interests can easily be in direct conflict with one another (Hutchinson and Mitchell, 2019; Sonboli et al., 2021; Ekstrand et al., 2022; Smith et al., 2023b).
 A system that prioritizes promoting content from historically marginalized providers might inadvertently serve users content that is misaligned with their previously stated interests (Smith et al., 2023b).
-And as Smith et al.
-(2024) point out, such a system can even have potentially negative effects for providers by exposing them to harassment from users who are outside of their preferred target audience.
+In fact, recent user studies have found that users often prefer gender-biased recommendations over ‘debiased’ ones, even when they are made aware that bias is present (Wang et al., 2023; Krause et al., 2025).
+As Smith et al.
+(2024) point out, a ‘debiased’ system can even have potentially negative effects for providers by exposing them to harassment from users who are outside of their preferred target audience.
 Complicating things further, a single platform may contain multiple types of providers, such as social media content creators and paid advertisers, whose interests may conflict as well.
-Despite the fact that RS are a commonly used technology, much ML fairness research has focused on classification problems (Barocas and Selbst, 2016; Chouldechova, 2017; Kleinberg et al., 2017; Dwork et al., 2012; Feldman et al., 2015; Zafar et al., 2017a; Corbett-Davies et al., 2017, 2023; Friedler et al., 2019; Kilbertus et al., 2017; Hardt et al., 2016b; Pleiss et al., 2017; Kamiran and Calders, 2009; Buolamwini and Gebru, 2018; Obermeyer et al., 2019; Friedler et al., 2021; Lee et al., 2024; Grgic-Hlaca et al., 2016; Hardt et al., 2016a; Sharifi-Malvajerdi et al., 2019; Zafar et al., 2017b; Borgs et al., 2019; Tal, 2023)—meaning that many existing fairness objectives and metrics cannot be directly applied to RS.
-Large-scale RS also raise unique challenges in practice related to sparse user-content interaction data, because users may only interact with a tiny subset of overall content (Li et al., 2021a; Kearns et al., 2019; Beutel et al., 2019b, a, 2018; Lum et al., 2022).
-Finally, as RS use user-content interaction data generated from the recommendation process to inform their future results, they risk feedback loops that amplify disparities (Keller, 2021).
-Our study contributes perspectives on implementing ML fairness in practice that are specific to RS, taking these unique challenges into account.
+Even within a stakeholder group, how fairness should be defined is not straightforward.
+For example, when considering providers in a video-sharing platform, while one approach to fairness might involve rewarding providers for producing high-quality content by using straightforward viewer metrics for ranking videos, an alternative approach might incorporate randomness in search rankings to counteract the “rich-get-richer” effect (Chen et al., 2023) wherein top-ranked videos get disproportionately more future views.
+The massive scale of RS also raises unique challenges in practice.
+As Anthis et al.
+(2025) argue, technologies that are used by large populations across many contexts, like RS, simply cannot be made fair for every context in which they operate and every population subgroup with which they interact.
+Scale also causes technical challenges related to sparse user-content interaction data, because users may only interact with a tiny subset of overall content (Li et al., 2021a; Kearns et al., 2019; Beutel et al., 2019b, a, 2018; Lum et al., 2022).
+Finally, as RS use user-content interaction data generated from the recommendation process to inform their future results, they risk feedback loops that amplify disparities (Keller, 2021; Li et al., 2025).
+Feedback loops themselves are multi-faceted: users can provide implicit feedback (e.g., quickly scrolling past a video) or explicit feedback (e.g., requesting not to see a video again) (Oard et al., 1998; Li et al., 2025).
+Users provide explicit feedback less consistently than they provide implicit feedback (Lustig et al., 2022); however, implicit feedback is harder to interpret and is often itself shaped by users’ folk theories of how the RS they interact with work (Eslami et al., 2016; DeVito et al., 2017; Xiao et al., 2025).
+The choices by RS practitioners about what feedback to include—and the choices by users about how to convey and organize their feedback—can thus have major impacts on fairness in RS.
+Our study builds on the fair ML literature by examining how inherent fairness tensions are approached in practice and the extent to which academic definitions of fairness guide industry approaches, taking the unique challenges of RS fairness into account.
 2.2. Fairness in Practice
+Bringing About Fairness Efforts
 In practice, fairness efforts do not occur in isolation.
@@
 Acquiring organizational support to conduct responsible AI initiatives often requires linking fairness to explicit business or product goals (Scheuerman, 2024; Ali et al., 2023; Garcia-Gathright et al., 2018; Rakova et al., 2021; Veale et al., 2018; Wang et al., 2024a).
-Legal teams can also play a role in bringing about responsible AI efforts, although they often focus on privacy or liability rather than proactive fairness (Veale and Binns, 2017).
+Regulation can also play a role in bringing about responsible AI efforts.
 However, regulations often vary across legal jurisdictions, complicating fairness efforts for multinational platforms that use RS.
-For example, RS in the EU are subject to the Digital Services Act (DSA), which authorizes authorities to enforce yearly assessments of online platforms regarding fairness issues, illegal content, and disruption to human rights (Madiega, 2020).
-The DSA does not apply to RS in the US (Fiordalisi, 2025); however, individual states have created their own fairness regulations, such as California’s 2022 state law AB 587, [^2] https://openstates.org/ca/bills/20212022/AB587/ which aims to provide transparency about large platform recommendations to avoid disinformation and hate speech.
-When fairness standards are not mandated by law, legal teams may actually hinder fairness efforts due to concerns that they could add to organizational liability (Costanza-Chock et al., 2022; Wright et al., 2024).
-If institutional support is lacking, employees may choose to “individualize risk” (Ali et al., 2023) by taking personal responsibility for incorporating fairness into their workflows.
-Prior work has repeatedly shown that these efforts can be challenging, as practitioners struggle to reconcile incompatible fairness metrics (Madaio et al., 2022; Moharana et al., 2025; Smith et al., 2025), work with messy real-world data (Cramer et al., 2019; Holstein et al., 2019; Rieke et al., 2022; Guerdan et al., 2025; Passi and Barocas, 2019; Sambasivan et al., 2021), and adapt general-purpose academic tools and metrics to their specific real-world needs (Deng et al., 2022; Harvey et al., 2025, 2024; Lee and Singh, 2021; Balayn et al., 2023; Kaur et al., 2020; Richardson et al., 2021; Ojewale et al., 2025).
-Many of these challenges are exacerbated by organizational challenges, such as communicating and collaborating across different teams and within cross-functional teams (Deng et al., 2023b).
-Researchers including Madaio et al.
-(2020), Deng et al.
-(2023a), and Costanza-Chock et al.
-(2022) have repeatedly found that a particular pain point when attempting to improve algorithmic fairness in practice is stakeholder involvement: practitioners struggle to involve the perspectives of end users in fairness assessments and improvements.
+For example, RS in the EU are subject to the Digital Services Act (DSA), which authorizes yearly assessments of online platforms regarding fairness issues, illegal content, and disruption to human rights (Madiega, 2020).
+The DSA does not apply to RS in the US (Fiordalisi, 2025); however, individual states have created their own fairness regulations, such as California’s 2022 state law AB 587, [^3] https://openstates.org/ca/bills/20212022/AB587/ which aims to provide transparency about large social media RS to avoid disinformation and hate speech.
+When fairness standards are not mandated by law, organizations may avoid assessing fairness altogether due to concerns that uncovering unfairness could add to organizational liability (Costanza-Chock et al., 2022; Wright et al., 2024).
+If institutional support is lacking, employees may choose to “individualize risk” (Ali et al., 2023) by taking personal responsibility for incorporating fairness into their workflows (Ganesh et al., 2025).
+Operationalizing Fairness
+However efforts are initiated, doing fairness in practice requires adhering to a specific operationalization of fairness.
+For example, documentation efforts like datasheets for datasets (Gebru et al., 2021), model cards for model reporting (Mitchell et al., 2019), and FactSheets (Arnold et al., 2019) call on practitioners to publicly share standardized reports describing the intended use, disaggregated performance, and other fairness-related characteristics of their data and models.
+Similarly, fairness checklists, which are inspired by safety checklists used in medicine and aviation, are intended to encourage practitioners to consider fairness issues and mitigate potential sources of unfairness throughout the design and development of technical systems (Madaio et al., 2020).
+As a more technical operationalization, much fair ML research focuses on developing metrics and toolkits that practitioners can use to automatically measure and improve fairness in decision-making systems (Ojewale et al., 2025; Wong et al., 2023).
+Fairness Challenges in Practice
+Prior work has repeatedly shown that efforts to operationalize fairness within organizations are deeply challenging.
+Practitioners struggle to reconcile incompatible fairness metrics (Madaio et al., 2022; Moharana et al., 2025; Smith et al., 2025; Koenecke et al., 2023), work with messy real-world data (Cramer et al., 2019; Holstein et al., 2019; Rieke et al., 2022; Guerdan et al., 2025; Passi and Barocas, 2019; Sambasivan et al., 2021; Pang et al., 2023), and adapt general-purpose academic tools and metrics to their specific real-world needs (Deng et al., 2022; Harvey et al., 2025, 2024; Lee and Singh, 2021; Balayn et al., 2023; Kaur et al., 2020; Richardson et al., 2021; Ojewale et al., 2025; Madaio et al., 2024; Morley et al., 2020; Corpus et al., 2025).
+Challenges are exacerbated by organizational barriers to communicating and collaborating across different teams and within cross-functional teams (Deng et al., 2023b), or to incorporating the perspectives of multiple stakeholders, including end users, into fairness assessments and improvements (Deng et al., 2023a; Costanza-Chock et al., 2022; Madaio et al., 2020).
 With the notable exception of Smith et al.
@@
 Most similar to our work, Smith et al.
-have also conducted interview studies with RS practitioners themselves: employees of an anonymous “commercial consumer-provider recommender system” (Smith et al., 2023a) and of a microlending platform (Smith et al., 2023b).
+have conducted interview studies with RS practitioners themselves: employees of an anonymous “commercial consumer-provider recommender system” (Smith et al., 2023a) and of a microlending platform (Smith et al., 2023b).
 Both studies explored challenges related to selecting fairness metrics in the context of RS, again identifying competing stakeholder interests as a key challenge (Smith et al., 2023b, a).
 In contrast, we interview technical RS practitioners across multiple big tech companies, and look beyond fairness metrics to how fairness is incorporated into RS workflows more broadly.
-In summary, we build on the extensive body of prior work by focusing explicitly on the RS workflow, with a particular spotlight not only on how RS practitioners across multiple big tech companies face technical challenges in the face of RS-specific complexity, but also how they navigate organizational challenges stemming from cross-team interactions with legal, data, and fairness practitioners.
+In summary, we build on the extensive body of prior work by focusing explicitly on the RS workflow, with a spotlight on how RS practitioners across multiple big tech companies navigate not only technical challenges stemming from RS-specific complexity, but also organizational challenges stemming from cross-team interactions with legal, data, and fairness practitioners.
 3. Methods
@@
 We interviewed RS practitioners from a variety of large companies that deploy highly visible RS.
+Our interviews were intended to answer two key research questions:
+  • RQ1:
+    What do the technical and fairness workflows of an RS practitioner look like, and how do they intersect?
+  • RQ2:
+    What challenges do RS practitioners face when incorporating fairness considerations into their technical workflows?
 Recruitment and Participants
@@
 Practitioners held titles such as ‘Machine Learning Engineer,’ ‘Machine Learning Researcher,’ or ‘Scientist,’ and were actively involved in building RS.
-Each participant worked for a company that serves over one billion customers annually.
+We specifically focused on technical practitioners because we were interested in understanding how fairness considerations are integrated into the work of building RS and how practitioners handle dual technical and fairness responsibilities.
+(We discuss the implications of this decision in more detail in §8.1.) Each participant worked for a company that serves over one billion customers annually.
 Further details on participants are available in Table 1.
@@
 P11 12
-Table 1. Participant Background Information.:
+Table 1. Participant Background Information.
+A table describing participant background.
+There are 11 participants (P1-P11).
+All participants held roles that involved engineering responsibilities, and all but P5 and P7 held roles that involved research responsibilities.
+Participants had varying years of experience: 3, 4, 4, 5, 6, 5, 4, 5, 7, 3, and 12, respectively.
 Interviews
 Interviews ranged between 45 minutes and one hour.
-We began by asking ice-breaking questions regarding the practitioner’s background, including their years of experience, employer, and the RS that they built or worked with.
 Practitioners were instructed to be as detailed as possible but to avoid sharing any confidential information such as revenue or upcoming products.
-Next, we asked questions intended to zero in on practitioner workflows and approaches to fairness.
-We prepared a set of initial opening questions to encourage practitioners to share their general experiences and provide targets for follow-ups.
-We then proceeded to a set of follow-up questions about fairness metrics, practitioner time allocation, and cross-team collaboration based on earlier responses.
-Finally, we invited practitioners to comment on and rate their professional experiences in RS in terms of emotional support, work intensity, and stress level.
+We began by asking ice-breaking questions regarding practitioner background, including their years of experience, employer, and the RS that they built or worked with.
+Next, we asked questions intended to help us map practitioner workflows (RQ1).
+We began with open-ended questions (e.g., “could you describe a recent project’s workflow?”, “are there external teams that you interact with regularly?”) and asked more detailed follow-up questions if necessary (e.g., “how do you feel when interacting with the data team?”).
+We chose to begin with these open-ended, workflow-centric questions in order to develop an understanding of how central fairness work was to practitioners’ broad descriptions of their responsibilities before they were explicitly prompted to consider fairness.
+Next, we asked about participants’ professional experience with assessing and improving fairness in RS.
+These included workflow-related questions intended to address RQ1 (e.g., “what amount of time do you spend on fairness in your daily work?”) as well as questions related to potential challenges intended to address RQ2 (e.g., “can you describe a recent time that improving fairness was difficult or failed?”).
+We also asked about participants’ perceptions of their organizations’ approaches to fairness (e.g., “is fairness in your organization a team effort?”).
+We asked about these separately to try to disentagle participants’ technical considerations and challenges from organizational ones.
+Interview quotes included in this paper have been lightly edited for clarity.
 Our semi-structured interview guide is provided in Appendix Table 3.
@@
 We sought to understand how fairness is approached in practice by individuals and teams who develop and deploy large-scale RS.
-To do so, we analyzed responses on two levels: an affinity diagramming exercise to understand RS practitioner workflows, and a thematic analysis to understand the challenges faced by RS practitioners who seek to improve fairness in the systems that they build, deploy, and maintain.
-Affinity Diagramming.
-First, we conducted an affinity diagramming exercise to map practitioners’ workflows based on interview notes and transcripts, following Yang et al.
-(2016).
-We focused on identifying commonalities, and also flagged outlier cases in our notes for further examination.
-We identified key stages practitioners shared in building RS.
-We also identified the core set of collaborators and external teams with whom RS practitioners interacted in order to construct a flow model of communications (Wixon et al., 1990).
-Data were analyzed during collaborative meetings between authors, with interviewers summarizing their sessions and outlining notes and/or transcripts relevant to each session under inspection.
-We then collaboratively sketched and refined models describing both the general flow of information and data during model development as well as the specific fairness approaches undertaken by practitioners.
+To do so, we analyzed responses on two levels: first, we mapped RS practitioner workflows, and second, we conducted an inductive thematic analysis to understand the challenges faced by RS practitioners who seek to improve fairness in the systems that they build, deploy, and maintain.
+RQ1: Mapping Practitioner Workflows.
+First, we developed a unified mapping of practitioners’ workflows.
+As our goal was to recognize general patterns among practitioners, we focused on identifying commonalities and flagged outlier cases for discussion and further examination.
+Data were analyzed during collaborative meetings between authors, with interviewers summarizing their sessions and outlining transcripts and/or notes relevant to each session under inspection.
+As each transcript or set of notes was discussed, we identified key stages to building an RS that were discussed by practitioners.
+We also identified the core set of external teams with whom practitioners interacted.
+We then collaboratively sketched and refined workflow models describing both the general flow of data during model development as well as the specific fairness approaches undertaken by practitioners at each stage of the process.
 Findings are described in §4.
-Thematic Analysis.
-Second, we conducted a thematic analysis to identify core challenges faced by practitioners (Braun and Clarke, 2019, 2006).
-We first converted scripts and notes into items on a Miro web board for aggregation and analysis.
-Authors conducted two coding passes through all quotes and reflections, developing a large set of initial theme groups.
-In the second round, authors aggregated thematic groups, merging similar themes and recategorizing elements through negotiation.
+RQ2: Understanding Practitioner Challenges.
+Second, we conducted an inductive thematic analysis to identify core challenges faced by practitioners as they sought to incorporate fairness considerations into their technical workflow (Braun and Clarke, 2006).
+We first converted transcripts and interview notes into items on a Miro web board for aggregation and analysis.
+Authors conducted two coding passes through all quotes and notes, developing a large set of initial codes related to any mention of challenges that practitioners faced when trying to incorporate fairness into their technical RS workflows.
+Next, authors collated codes into initial themes.
+Then, authors reviewed and refined themes, aggregating thematic groups, merging similar themes, and recategorizing elements through negotiation.
 Lastly, through another open discussion, authors broke down themes that had a significantly larger number of notes into finer granularity and placed any remaining unlabeled items into categories.
 When finalizing results, authors summarized the high-level themes identified and pulled representative quotes from the grouped items.
-Findings about technical challenges faced by RS practitioners are described in §5 and findings about organizational challenges arising from RS practitioner interactions with other teams are described in §6.
+Findings about technical challenges faced by RS practitioners are described in §5, and findings about organizational challenges arising from RS practitioner interactions with other teams are described in §6.
 Positionality
 We are a team of researchers who work on academic and industry teams, and who collectively have expertise in machine learning, human-computer interaction, and algorithmic fairness.
-Our backgrounds have certainly informed our approach to this study.
+Our backgrounds have informed our approach to this study.
 Our experience on industry teams informed our understanding of the practical constraints that practitioners may face, which influenced the questions we asked in our semi-structured interviews.
 At the same time, our experience does not extend to working as RS practitioners within technology companies.
-While this positions us to critically assess RS workflows and challenges as outsiders, we do by definition lack ‘insider’ knowledge about what it is like to work on RS in practice.
+While this positions us to critically assess RS workflows and challenges as outsiders, we lack ‘insider’ knowledge about what it is like to work on RS in practice.
 Finally, we are also all end-users of RS systems.
-We believe that this dual vantage point as both researchers and users shaped our understanding of the potential real-world impacts of RS, which in turn informed our approach to thematic analysis.
+This dual vantage point as both researchers and users shaped our understanding of the potential real-world impacts of RS, which in turn informed our approach to thematic analysis.
 4. The RS Practitioner Workflow
@@
 We visualize RS practitioners’ interactions with the data team managing and monitoring production data access, the legal team auditing and approving specific data access requests, and the fairness (or responsible AI) team providing fairness expertise.
-Distinct colors—blue for technical (RS practitioners), yellow for legal, red for data, and green for fairness—represent team inputs.
-Dotted arrows indicate occasional, rather than consistent, participation by a team in certain stages.
-Offline phases are denoted as gray boxes, while online phases are in purple boxes.
-Green circles indicate phases in which fairness considerations are commonly incorporated.:
+Dotted borders indicate offline phases of the workflow, and solid borders indicate online phases.
+For each phase of the technical workflow (top row), there are associated phases within the fairness workflow (second row).
+In each phase, different participants interact (bottom row).
+A figure summarizing the RS practitioner technical and fairness workflows.
+The technical workflow includes the following offline phases: data acquisition and processing, model construction, and model validation; as well as the following online phases: model testing and model service.
+The fairness workflow co-occurs with the technical workflow.
+The fairness workflow consists of the prototyping, internal, and collaborative phases.
+The prototyping phase, which co-occurs with the offline data acquisition and processing and model construction phases, includes the technical team, the legal team, and the data team.
+In the protoyping phase, the following actions are taken: the technical team makes data requests, the legal team approves or rejects data requests based on regulatory or risk considerations, the data team facilitates data requests via API access, the technical team hypothesizes about sources of unfairness, and the technical team develops or adapts fairness metrics.
+The internal phase, which co-occurs with the offline model validation phase, includes the technical team.
+In the internal phase, the technical team validates fairness metrics, conducts iterative fairness testing, and manually searches for previously unknown instances of unfairness.
+The collaborative phase, which co-occurs with the online model testing and model service phases, includes the technical team, the fairness team, and end users.
+In the collaborative phase, the fairness team provides feedback based on internal guidelines; end users interact with the RS and provide implicit and explicit feedback; and the technical team manually searches for previously unknown instances of unfairness, determines which fairness and other issues to resolve, modifies fairness metrics, and conducts iterative fairness testing.
 In this section, we describe the RS practitioner workflow.
-We focus on how RS practitioners integrate fairness work into their labor, both internally (within technical teams) and in collaboration with legal, data, and fairness teams.
-These findings emerge out of the workflow-centric affinity diagramming exercise described in §3.
+We focus on how RS practitioners integrate fairness work internally (within technical teams) and in collaboration with legal, data, and fairness teams.
 4.1. An Overview of the RS Practitioner Workflow
-The RS practitioner workflow is summarized in Figure 1.
-Practitioners described workflows that evolved from an offline phase (gray boxes in Figure 1) into an online phase (purple boxes in Figure 1) throughout development.
-First, practitioners coordinate with the legal and data teams to receive access to necessary data and begin the data processing phase.
-The data team negotiates technical details such as API designs for collecting data and data cleaning, while the legal team reviews data access requests.
-Practitioners then spend much of their time on model construction and conducting rounds of offline model validation.
-All agreed that they invested more than half of their work time into these model-centric activities.
+The RS practitioner workflow is summarized in Figure 1. Practitioners described workflows that evolved from an offline into an online phase throughout development.
+Offline Development
+First, practitioners coordinate with the legal and data teams to receive access to necessary data and begin the data acquisition and processing phase.
+The data team negotiates technical details such as API designs for data collection as well as data cleaning, while the legal team reviews data access requests.
+Practitioners then spend much of their time on model construction and conducting rounds of model validation.
+All participants agreed that they invested more than half of their work time into these model-centric activities.
 For example, P5 reported, “This is the core of our daily work.
-I spent more than 60% time on model work.” Overall, the offline stages of the RS practitioner workflow align with the stages of general ML pipelines, e.g., as described by Suresh and Guttag (2021) and Black et al.
+I spend more than 60% [of my] time on model work.” Overall, the offline stages of the RS practitioner workflow align with the pre-deployment stages of general ML pipelines as described by Suresh and Guttag (2021) and Black et al.
 (2023), as well as with preferences towards “model work” over “data work” that have been previously documented by Sambasivan et al.
 (2021).
-After offline development, practitioners ship their model and move to an online testing phase, where real-world user behavior data and newly generated user-content interaction data are collected as users engage with recommendations.
-A fairness team, which all practitioners reported exists in their organizations under various names (such as ‘Responsible AI’), provides feedback to RS practitioners at this stage (see §4.2).
-Simultaneously, practitioners monitor the model’s performance metrics during the online model service phase, and might revisit the data pre-processing phase to change their training data (e.g., performing feature transformations or asking for more features).
+Online Development
+After offline development, practitioners ship their model and move to a model testing phase, where real-world user behavior data and newly generated user-content interaction data are collected as users engage with recommendations.
+A fairness team, which all participants reported exists in their organizations under various names (such as ‘responsible AI’), provides feedback to RS practitioners at this stage (see §4.2).
+Simultaneously, practitioners monitor model performance metrics during the model service phase, and might revisit the data acquisition and pre-processing phase to change their training data (e.g., performing feature transformations or asking for more features).
+These online stages of the RS practitioner workflow align with the post-deployment stages of general ML pipelines (Suresh and Guttag, 2021; Black et al., 2023).
+However, while prior work has identified only a small amount of ML fairness research focused on post-deployment efforts (Black et al., 2023), participants described the online phases as a critical—and long-running—part of their workflow.
+As P2 put it, “We are in [a] loop [where we keep] validating, online testing and improving the model.
 4.2. How the RS Practitioner Workflow Incorporates Fairness
-Fairness assessments and improvements occurred at several points throughout the RS workflow, corresponding to shifts to new development stages (green circles in Figure 1).
+Fairness assessments and improvements occurred at several points throughout the RS workflow, often corresponding to shifts to new development stages.
 These transition points were convenient locations to incorporate fairness because they often included interactions between different teams.
@@
 It consists of three major phases, as depicted in Figure 1: the prototyping phase, the internal phase, and the collaborative phase.
-The prototyping phase co-occurs with the data processing phase of the RS workflow.
-In the data processing phase, RS practitioners interacted with legal and data teams to obtain access to training data.
-Concurrently, practitioners started the prototyping phase by hypothesizing about the sources of unfairness that might exist in the training data.
-Practitioners based hypotheses on their prior experience with RS, and did not generally cite academic research as a source for their hypotheses.
-Likewise, practitioners reported reusing fairness metrics that they had previously developed to assess training data in this stage.
-Later, during the offline model validation phase of the RS workflow, practitioners moved to the internal phase of the fairness process, wherein they validated the metrics developed during the prototyping phase and manually searched for unexpected results.
-Practitioners expressed that there are several rounds of iteration at this stage to find as many instances of potential unfairness as possible prior to model deployment.
-Figure 2. Recommender System Practitioner Fairness Process. Technical teams incorporate fairness into their workflows in collaboration with legal, data, and fairness teams. :
-After the model is shipped for online testing, the fairness process enters the collaborative phase.
+Prototyping Phase
+The prototyping phase co-occurs with the data acquisition and processing phase of the RS workflow.
+In the data acquisition and processing phase, technical RS practitioners interact with legal and data teams to obtain access to training data.
+Here, procedural fairness considerations are raised by the first time as the legal team communicates to the technical team which features (e.g., sensitive attributes) should not be used in building RS due to regulatory requirements or company policy.
+Unlike prior work, which has often found AI practitioners to be at odds with legal teams who see measuring unfairness as a potential legal liability (Costanza-Chock et al., 2022; Wright et al., 2024), we found that participants had overwhelmingly positive opinions of the legal team (for example, P2 saw the legal team as “help[ing] us comply well with our company’s vision and values.”).
+Concurrently, practitioners spend time during the prototyping phase hypothesizing about the sources of unfairness that might exist in the training data and developing metrics intended to examine those hypotheses.
+Participants based hypotheses on their prior experiences with RS, and did not generally cite academic research as a source.
+P9 stated that even if they were aware of relevant academic research on fairness considerations related to their use case, they still “need to come up with a clear formula and implement it by ourselves—it is really domain-specific.” Instead, participants reported reusing fairness metrics that they had previously developed to assess training data in this stage and emphasized that their initial metrics were “intuition-based” (P3).
+A similar reliance on bespoke as opposed to standardized fairness metrics and approaches has been noted in prior work on NLP practitioners (Harvey et al., 2025) as well as AI practitioners generally (Madaio et al., 2022, 2024).
+Internal Phase
+During the offline model validation phase of the RS workflow, practitioners moved to the internal phase of the fairness workflow, wherein they validated metrics developed during the prototyping phase and manually searched for unexpected results.
+Participants expressed that they undertake several rounds of iteration in this phase to find as many instances of potential unfairness as possible prior to model deployment.
+In addition, participants emphasized the uncertain nature of the internal phase.
+According to P1, “It is really hard to predict users’ behaviors—a lot of uncertainties.” P2 described the internal phase bluntly: “We have to look carefully to find those hidden biases.
+It is not fun.” These comments point to a practical challenge that has not been widely discussed in prior work: because RS are an inherently collaborative technology, testing them requires a model of user behavior.
+Thus, relying on historical data as a test set, as is commonly done in prediction tasks, is not sufficient in an RS setting.
+Collaborative Phase
+After the model is shipped for online testing, the fairness workflow enters the collaborative phase.
 This phase is typically the first to include the fairness team, as well as the first phase to include data from real-world users of the RS.
 In the collaborative phase, the fairness team offers feedback based on their internal guidelines.
-Most practitioners reported that they had discretion as to whether to integrate feedback from the fairness team.
-Because practitioners felt that they had limited time and resources available, they sometimes attempted to weigh the relative impact of potential fairness issues against the work required to resolve them.
-As the owners of an RS within their organizations, practitioners reported numerous ongoing job requirements, such as monitoring the model’s online performance.
-Practitioners primarily reported being concerned about factors such as uptime, and did not consider fairness to reach a similar level of disruptive emergency necessitating urgent correction.
-When RS practitioners were able to address fairness issues in deployed models, they often had to integrate user feedback and data to seek out ‘hidden’ types of unfairness (i.e., those that they had not anticipated in the prototyping stage).
+Most participants reported that they had discretion as to whether to integrate feedback from the fairness team.
+Because participants felt that they had limited time and resources available, they sometimes attempted to weigh the relative impact of potential fairness issues against the work required to resolve them:
+P1: If [the fairness team’s] feedback is about removing one particular feature, normally we will simply do it.
+However, if it requires a lot of work, we probably need to have another broader discussion.
+As the owners of an RS within their organizations, participants reported numerous ongoing job requirements, such as monitoring online model performance.
+Participants primarily reported being concerned about factors such as uptime, and did not consider fairness to reach a similar level of disruptive emergency necessitating urgent correction.
+These types of competing demands on practitioners’ time have been repeatedly noted by prior research (Wang et al., 2024a; Madaio et al., 2022).
+However, the iterative, manual, and user behavior-dependent aspects of fairness measurement are more specific to RS.
+When participants were able to address fairness issues in deployed models, they often had to integrate user feedback and data to seek out “hidden” (P2, P7) types of unfairness (i.e., those that they had not anticipated in the prototyping stage).
 This process is highly iterative and requires much manual analysis.
-Practitioners mentioned that, in a typical project, they eventually compose over five new fairness metrics tailored to the specific system and context, even though they typically started with pre-existing metrics from prior projects.
+Participants reported that, in a typical project, they eventually composed over five new fairness metrics tailored to the specific system and context, even though they typically started with pre-existing metrics from prior projects.
 5. Technical Challenges Arising Within RS Practitioner Teams When Incorporating Fairness
@@
 (Within RS Teams)
-Defining Fairness Defining fairness in specific RS contexts, especially those not subject to pre-existing rules or regulation (11/11) Adapting pre-existing fairness metrics for RS (7/11), and managing the internal documentation involved in those adaptations (8/11)
-Considering Multiple Types of Fairness Across Many Groups Managing (conflicting) multi-stakeholder interests (6/11) Prioritizing between different (non-conflicting) types of fairness (8/11) Handling high numbers of user groups (5/11)
-Dynamic Fairness Considerations Developing ad hoc metrics to detect unexpected sources of bias, especially when managing real-world user interaction and feedback loops (10/11)
-Organizational (Cross-Team) Lack of Time Finding time to work on fairness within technical teams (9/11) and in cross-team collaborations (6/11)
-Cross-Team Communication Communicating and collaborating effectively with external teams (7/11)
-Table 2.
-Challenges Faced by RS Practitioners when Incorporating Fairness into their Workflows.
-Challenges occur within RS teams as well as across teams and at the organization level.
-The number of practitioners who reported experiencing each type of challenge is shown in parentheses.:
-In this section, we describe the key technical challenges faced by RS practitioners who seek to incorporate fairness considerations into their workflows.
-These findings emerge out of the challenge-centric thematic analysis described in §3.
-Table 2 details the technical challenges most commonly reported by practitioners.
+Defining Fairness Defining fairness in specific RS contexts, especially those not subject to existing rules or regulations Adapting pre-existing fairness metrics for RS, and managing the internal documentation involved in those adaptations
+Considering Multiple Types of Fairness Across Many Groups Managing (conflicting) multi-stakeholder interests Prioritizing between different (non-conflicting) types of fairness Handling high numbers of user groups
+Dynamic Fairness Considerations Developing ad hoc metrics to detect unexpected sources of bias Managing real-world user interaction and feedback loops
+Organizational (Cross-Team) Lack of Time Finding time to work on fairness within technical teams and in cross-team collaborations
+Cross-Team Communication Communicating and collaborating effectively with external teams
+Table 2. Challenges Faced by RS Practitioners when Incorporating Fairness into their Workflows. Challenges occur within RS teams as well as across teams and at the organization level.
+A table summarizing challenges faced by RS practitioners when incorporating fairness into their workflows.
+Technical challenges (within RS teams) include defining fairness, considering multiple types of fairness across many groups, and dynamic fairness considerations.
+Defining fairness includes defining fairness in specific RS contexts, especially those not subject to pre-existing rules or regulation, adapting pre-existing fairness metrics for RS, and managing the internal documentation involved in those adaptations.
+Considering multiple types of fairness across many groups includes managing (conflicting) multi-stakeholder interests, prioritizing between different (non-conflicting) types of fairness, and handling high numbers of user groups.
+Dynamic fairness considerations include developing ad hoc metrics to detect unexpected sources of bias, especially when managing real-world user interaction and feedback loops.
+Organizational challenges (cross-team) include lack of time (finding time to work on fairness within technical teams and in cross-team collaborations) and cross-team communication (communicating and collaborating effectively with external teams).
+In this section, we describe the key technical challenges faced by RS practitioners who seek to incorporate fairness considerations into their workflows (summarized in Table 2).
+Overall, we find that defining fairness in the first place is a challenge, echoing findings from prior interview studies conducted on AI and ML practitioners (Madaio et al., 2022).
+We also find challenges that are more unique to RS, such as that RS practitioners must consider multiple types of fairness across many different stakeholder groups.
+RS practitioners must not only balance conflicting stakeholder interests, but also prioritize between a large number of different non-conflicting types of fairness.
+In addition, RS practitioners face unique challenges related to assessing and improving fairness in dynamic settings.
+Because RS are collaborative systems that are shaped by user behavior, pre-deployment fairness assessments are especially challenging.
 5.1. RS Practitioners Struggle to Define Fairness in the Context of Recommender Systems
-Defining Fairness.
-Practitioners attempted to assess and improve fairness within two general categories.
+Defining Fairness
+RS practitioners attempt to assess and improve fairness within two general categories.
 The first category is fairness as mandated by existing rules or regulations.
-For example, an RS that discriminated against job candidates from particular races would be considered unfair under US anti-discrimination law.
-The second category, encompassing all contexts not covered by existing rules or regulations, is perceived as more difficult and nuanced.
-One common frustration was that practitioners only have a vague goal (for example, not exposing content creators to accounts that might harass them), but are not armed with specific tools in their organization to reach the goal—or even a standardized method with which to approach the problem.
-Adapting Pre-Existing Fairness Metrics and Navigating Documentation Debt.
-While approaches for assessing and improving fairness have been proposed by fair ML literature (Chen et al., 2023), practitioners found them hard to directly adapt and ill-suited for their existing workflows.
-For example, P9 reported, “Even though we have a list of categories [over which to assess fairness] as you mentioned in research papers, we need to come up with a clear formula and implement it ourselves.
-It is really domain-specific.” Practitioners also struggled to build and re-use institutional knowledge about fairness in RS.
-This was due to the fact that RS practitioners faced documentation debt (Bender et al., 2021): prior fairness-related work at their organizations was not typically well-documented for future use.
-Half of practitioners interviewed therefore faced difficulties in re-using internal metrics that had been previously developed for fairness purposes.
-P7 shared, “For sure we can find their code, and comments, but without some context archived, for example, some documents, we are not sure whether we can use [them] directly.” Practitioners also felt ill-equipped to document their own fairness processes.
-As P3 stated, “Documentation is a good idea…however, I am not sure how specific of my debiasing should be included in the documentation.”
+For example, an RS that discriminates against job candidates of particular races would be considered unfair under US anti-discrimination law.
+Even when participants had a legally mandated fairness goal in mind, they reported struggling with how to achieve it.
+Participants felt they were not armed with specific tools or standardized methods to reach even well-specified fairness goals.
+This echoes Smith et al.
+(2023a), who found that as RS practitioners tried to select appropriate fairness metrics, they struggled with not only normative questions like which stakeholders should be prioritized, but also practical ones like whether to measure fairness across an entire distribution of recommendations or only across a top-k set of rankings.
+Participants perceived the second category, encompassing all contexts not covered by existing rules or regulations, as more difficult and nuanced.
+This category includes cases where attributes that are legally protected in some contexts are not legally protected in others.
+For example, although racial discrimination is illegal in employment RS, it is not illegal in dating RS (Hutson et al., 2018).
+Thus, it is not clear how consistently to apply the same fairness metrics and techniques across contexts.
+This category also includes fairness across attributes that are not legally protected in any context, such as the popularity of a provider’s content.
+These attributes frequently conflict and can compound to the point of being overwhelming.
+For example, P7 stated that, “Each module of my recommender system requires different debiasing,” with goals ranging from avoiding a “rich gets richer” feedback loop to providing recommendations of equal quality to different groups of end users.
+This question of what to do in the absence of legal requirements is one that industry AI teams consistently grapple with (Quiñonero Candela et al., 2023).
+Adapting Pre-Existing Fairness Metrics and Navigating Documentation Debt
+While approaches for assessing and improving fairness in RS have been proposed by researchers (see §2.1), participants found them hard to directly adapt and ill-suited for their existing workflows.
+In some cases, participants simply did not have the time to find existing metrics:
+P4: I didn’t find very useful literature but honestly, I didn’t have much time to do a serious literature review and read papers either.
+Participants also struggled to build and re-use institutional knowledge about fairness in RS.
+This was due to the fact that they faced documentation debt (Bender et al., 2021): prior fairness-related work at their organizations was not typically well-documented for future use.
+Half of practitioners interviewed therefore faced difficulties in re-using internal metrics that had been previously developed for fairness purposes:
+P7: For sure we can find [old] code and comments, but without some context archived, some documents, we are not sure whether we can use [them] directly.
+Participants also felt ill-equipped to document their own fairness processes.
+As P3 stated, “Documentation is a good idea…however, I am not sure how specific [I should be].” These challenges strongly echo those surfaced by prior work on NLP, ML, and AI practitioners (Harvey et al., 2025; Holstein et al., 2019; Ojewale et al., 2025; Lee and Singh, 2021).
 5.2. RS Practitioners Must Account for Multiple Types of Fairness Across Many Groups
 Managing Conflicting Multi-Stakeholder Interests
-Prior research emphasizes the dual role of recommender systems in simultaneously connecting and serving multiple stakeholders, including providers and end-users (Burke, 2017; Li et al., 2021b).
+RS simultaneously connect and serve multiple stakeholders, including providers and end-users (Burke, 2017; Li et al., 2021b; Smith et al., 2023a).
 This multifaceted interaction introduces complex challenges, especially in the context of fairness.
-The practitioners that we interviewed commonly made tradeoffs, often prioritizing one user group over others.
-However, deciding how to make these tradeoffs was a challenge, as highlighted by P9: “Tradeoffs of a different group of users, or accuracy of the recommendations are hard to be decided,” and P1, who pointed out a lack of a “compelling [Standard Operating Procedure] for those decisions.” Making tradeoffs, while common, may not always be necessary.
-Recent academic findings suggest that multi-objective optimization might not always adversely affect particular user groups (Ekstrand et al., 2022), indicating a disconnect between industry practices and academic research.
-Prioritizing Between Different (Non-Conflicting) Types of Fairness.
-Practitioners felt that there was more of a “what to prioritize” question in RS compared with traditional predictive models.
-In a traditional ML pipeline, fairness assessment focuses on one particular target (e.g., demographic parity).
-However, the practitioners that we interviewed expressed that they have to balance multiple fairness-related considerations.
-For example, one participant needed to consider demographic parity for job seekers in one region—and simultaneously, ensure that advertisers across multiple regions were adhering to different sets of fairness-related regulatory requirements.
+Participants commonly made tradeoffs, often prioritizing one user group over others.
+However, deciding how to make these tradeoffs was a challenge, as highlighted by P9, who felt that, “Tradeoffs [between] groups of users…are hard to [decide],” and P1, who pointed out a lack of a “compelling [standard operating procedure] for those decisions.” Deciding which stakeholders to prioritize is a known challenge across fair ML (Madaio et al., 2022).
+However, in the context of RS, it may not always be necessary.
+Recent academic findings suggest that multi-objective optimization might not always adversely affect particular user groups, indicating a disconnect between industry practices and academic research in the RS domain (Ekstrand et al., 2022).
+Prioritizing Between Different (Non-Conflicting) Types of Fairness
+Participants felt that there was more of a question of what to prioritize in the context of RS as compared to predictive models.
+In a predictive ML pipeline, fairness assessment typically focuses on one particular target (e.g., demographic parity).
+However, the practitioners that we interviewed expressed that they had to balance multiple fairness-related considerations.
+For example, one participant needed to consider demographic parity for job seekers in one region—and simultaneously ensure that advertisers across multiple regions were adhering to different sets of fairness-related regulatory requirements.
 Importantly, these fairness considerations were not necessarily in conflict with one another.
-However, because practitioners had limited time and resources available, they nevertheless needed to choose certain fairness ideals to prioritize before turning to others.
-P3 pointed out: “Debiasing for multiple purposes is crucial in ensuring a fair and unbiased platform for all stakeholders.
-It’s not just about the debiasing of content creators based on region, gender, and religion…but also about addressing other factors that can impact the platform…For instance, we need to make sure that products…are supported fairly [e.g., that products that are artificially cheap because of government subsidies designed to gain market dominance are downranked on our RS platform].”
+However, because practitioners have limited time and resources available, they nevertheless need to choose certain fairness ideals to prioritize before turning to others:
+P3: Debiasing for multiple purposes is crucial in ensuring a fair and unbiased platform for all stakeholders.
+It’s not just about the debiasing of content creators based on region, gender, and religion…but also about addressing other factors that can impact the platform…For instance, we need to make sure that products…are supported fairly [e.g., that products that are artificially cheap because of government subsidies designed to gain market dominance are downranked on our RS platform].
+Because most work in fair ML focuses on prediction, in which different types of fairness are often incompatible with one another, prioritizing between different non-conflicting types of fairness represents a distinct challenge to doing fair ML in practice.
+While prior work has found that practitioners struggle to make value judgments about what to prioritize when faced with capacity constraints, that work tends to focus on the relationship between fairness and other, non-conflicting but not necessarily fairness-related, desiderata.
+For example, fairness efforts may be de-prioritized due to a focus on user privacy (Holstein et al., 2019) or an organizational emphasis on maximizing user engagement or product functionality (Rakova et al., 2021).
+Other work rightly notes that different fairness paradigms (e.g., procedural fairness and outcome-based fairness) do not necessarily conflict with one another and thus can be simultaneously satisfied by combining procedural fairness interventions (e.g., checklists) with technical ones intended to improve an outcome-based fairness metric (Madaio et al., 2020).
+In contrast, we focus on the idea that even within a single paradigm (e.g., outcome-based fairness), there are multiple non-conflicting fairness definitions that RS systems could simultaneously satisfy.
+Nevertheless, we find that even when a given practitioner is tasked with improving fairness (i.e., when fairness in general is not de-prioritized), that practitioner must essentially rank their planned fairness interventions (including both procedural and outcome-focused interventions) and prioritize accordingly.
 Handling High Numbers of User Groups
-A significant challenge identified by practitioners involves ensuring fairness across a wide range of groups within RS platforms.
+A significant challenge identified by participants involves ensuring fairness across a wide range of groups within RS platforms.
 Current metrics often fail to address the complexities arising from the large number of groups involved (and the ensuing intersectional groups).
-P2 described this challenge: “It’s hard to manage disparities between the numerous groups.
-With thousands of groups to consider, it becomes nearly impossible to ensure uniformity across all.” This sentiment was echoed by P9, who noted, “Most metrics are designed with the assumption of a few groups…They don’t cater to intra-group-level debiasing, making our adaptation efforts cumbersome.” Additionally, P7 highlighted the computational difficulties in handling sparse datasets resulting from subgroup division: “Most of the training data is too sparse, and it is hard for us to enforce the disparities between groups.
-I cannot look through thousands of groups to make sure that holds.”
+Multiple participants raised this challenge:
+P2: It’s hard to manage disparities between numerous groups. With thousands of groups to consider, it becomes nearly impossible to ensure uniformity across all.
+P7: Most of the training data is too sparse, and it is hard for us to enforce [fairness] between groups. I cannot look through thousands of groups to make sure [fairness] holds.
+Participants also struggled to find metrics that went beyond group fairness.
+For example, P9 noted that “Most metrics are designed with the assumption of a few groups…They don’t cater to intra-group-level debiasing, making our adaptation efforts cumbersome.” This challenge is exacerbated not only by the large scale of RS platforms, but also by the previously mentioned challenges of determining how to define and prioritize fairness metrics that are not required by law.
+In predictive ML settings, fairness assessments often take cues from anti-discrimination law by assessing differences in outcomes across legally protected demographic groups.
+In contrast, the RS practitioners we interviewed appear to rely less on anti-discrimination law as a blueprint for fairness assessments.
 5.3. Fairness Considerations are Complicated by Dynamic RS Environments
-Developing Ad Hoc Metrics to Detect Unexpected Instances of Unfairness.
+Developing Ad Hoc Metrics to Detect Unexpected Instances of Unfairness
 Practitioners engaged in building RS often create bespoke metrics in collaboration with their teams.
-P2 and P3 highlighted the personalized nature of these metrics, with P2 noting, “We proposed our own definitions of bias,” and P3 adding, “Most of the metrics we developed are intuition-based.” Such approaches underscore the reliance on practitioner expertise and intuition in assessing and improving fairness, as echoed by P6: “In general, I would like to write some hand-crafted codes to validate my hypothesis.
-Experience is a key to debiasing sometimes,” despite the availability of internal fairness tools.
-Nevertheless, practitioners acknowledged that their initial metrics often fell short in detecting unfairness, and they therefore engaged in iterative refinement of those metrics based on user feedback:
-[P1]: It is not an easy job…so we need to start from somewhere. After doing some debiasing based on your experience, we look forward to getting feedback from real-world users to further debiasing.
-[P7]: It sounds a bit contradictory. We have a model first and keep debiasing the model based on our user’s new data. It although risks bringing in some confirmatory bias, it works for now.
+Such approaches underscore the reliance on practitioner expertise and intuition in assessing and improving fairness, as highlighted by P6: “In general, I would like to write hand-crafted code to validate my hypothesis—experience is key to debiasing.” Nevertheless, participants acknowledged that their initial metrics often fell short in detecting unfairness, and they therefore engaged in iterative refinement of those metrics based on user feedback:
+P1: It is not an easy job…so we need to start from somewhere.
+After doing some debiasing based on [our] experience, we look forward to getting feedback from real-world users…We do everything that we can do.
+However, it is not rare we detect some other bias after the model is [deployed].
+P7: It sounds a bit contradictory. We have a model first and keep debiasing the model based on our users’ new data. Although [it] risks bringing in some confirmatory bias, it works for now.
+Again, this highlights a challenge specific to measuring and improving fairness in RS: offline testing is inherently limited because it requires detailed predictions of how users will interact with deployed RS.
+This means that RS practitioners must either develop models of predicted user behavior or must wait until systems are deployed for comprehensive testing.
+This stands in stark contrast to most predictive ML systems, which are often testable using historical data, and even to generative AI systems, which may be tested through red teaming.
 Managing Real-World User Interactions and Feedback Loops
-Practitioners also shed light on the challenges of assessing and improving fairness in online (deployed), as compared to offline, RS.
-This challenge is particularly acute in RS given the volume of data, the subtleties of human interactions, and the potential for user feedback loops to amplift latent sources of unfairness.
-P1 and P2 emphasized the difficulty in uncovering unexpected fairness issues in deployed systems, with P1 stating, “As the bias definitions are so different, and there are so many different user groups, those biases are hiding well behind the recommender systems that we built,” and P2 reflecting, “Most of the time, we have to look carefully to find those hidden biases.
-It is not fun.” Despite finding academic metrics less directly applicable (see §5.1), practitioners acknowledged the value of guidelines from research in providing a foundation and confidence in their decision-making processes.
-For example, P7 stated, “Though sometimes [checklists] are not directly related, I can still check those and they help me think through my reasoning.
-It made me somehow more confident in my decision.”
+Practitioners also face the challenge of assessing and improving fairness in deployed (as opposed to pre-deployment) RS.
+This challenge is particularly acute in RS given the volume of data, the subtleties of human interactions, and the potential for user feedback loops to amplify latent sources of unfairness.
+P1 emphasized the difficulty in uncovering unexpected fairness issues in deployed systems, stating, “As the bias definitions are so different, and there are so many different user groups, those biases are hiding behind the systems that we built.” Despite finding academic metrics less directly applicable (see §5.1), participants acknowledged the value of research in providing a foundation for their decision-making processes.
+For example, P7 stated, “Though sometimes [checklists] are not directly related, I can still check [them] and they help me think through my reasoning.
+It made me more confident in my decision.” At the same time, however, participants did not feel comfortable using pre-defined unfairness mitigation techniques (i.e., from academic research) on deployed systems.
+As P2 put it, “There is unknown risk of using those approaches.
+If the model performance drops a lot, it is a disaster.” This challenge is exacerbated by the fact that there is a comparatively small amount of fair ML research focused on post-deployment fairness improvements (Black et al., 2023).
 6. Organizational Challenges Arising Across Teams When Incorporating Fairness
-In this section, we describe the key organizational challenges faced by RS practitioners who seek to incorporate fairness considerations into their workflows.
-As with the technical challenges outlined in §5, these findings emerge out of the challenge-centric thematic analysis described in §3.
-Table 2 details the organizational challenges most commonly reported by practitioners.
+In this section, we describe the key organizational challenges faced by RS practitioners who seek to incorporate fairness considerations into their workflows (summarized in Table 2).
+The challenges we identify—namely, a lack of time to engage in fairness work and poor cross-team communication—echo those identified by prior interview studies conducted on AI and ML practitioners (Harvey et al., 2025; Deng et al., 2023b; Rakova et al., 2021; Balayn et al., 2023).
+However, we also identify a notable exception to the challenge of cross-team communication: while the RS practitioners we interviewed struggled to communicate with fairness teams, they overwhelmingly described clear and productive communications with legal teams.
+Thus, in this section, we contrast between the two to offer new insight into why RS practitioners struggle to develop a “fairness lingua franca.”
 6.1. RS Practitioners Lack Time to Engage in Collaborative Fairness Work
-Practitioners universally described their workflows as highly iterative and time-sensitive.
-As a result, most felt they had a very limited amount of time to focus on fairness, either internally or in collaborations with the fairness team.
-Nine reported that they spent roughly 5% percent of their work time on fairness-related tasks, and two said they spent up to 10% of their time.
-As P7 put it, “The negative effect spurred from the platform I worked for before is not so ‘bad.’ We don’t have the incentive to invest half of our time…to comprehensively debias.” When prompted to reflect on this allocation, most practitioners indicated that, while allocating more time for fairness-related tasks could be helpful, they were neutral about doing so.
-P1 reported, “Even though I wish I could have more time on this, I understand why this proportional[ly] makes sense.
-From the perspective of our org, prioritizing efforts on improving ‘important’ metrics brings more impact.” In other words, fairness issues did not rise to the level of a disruptive emergency, and thus were not prioritzed by management in evaluating project outcomes.
+Participants universally described their workflows as highly iterative and time-sensitive.
+As a result, most felt they had a very limited amount of time to focus on fairness, either internally or in collaborations with the fairness team: all reported spending 10% or less of their time on fairness-related tasks.
+As P7 put it, “The negative effect spurred from the platform I worked for before is not so ‘bad.’ We don’t have the incentive to invest half of our time…to comprehensively debias.” When prompted to reflect on this allocation, most participants indicated that, while allocating more time for fairness-related tasks could be helpful, they were neutral about doing so:
+P1: Even though I wish I could have more time on this, I understand why this proportionally makes sense.
+From the perspective of our [organization], prioritizing efforts on improving ‘important’ metrics brings more impact.
+In other words, fairness issues did not rise to the level of a disruptive emergency, and thus were not prioritzed by management in evaluating project outcomes.
+This challenge is not unique to RS, and has been identified by researchers time and time again as a particular barrier to doing fairness work in practice (Veale et al., 2018; Holstein et al., 2019; Madaio et al., 2020; Rakova et al., 2021; Madaio et al., 2022; Deng et al., 2023b; Balayn et al., 2023; Harvey et al., 2025).
 6.2. RS Practitioners Lack a ‘Fairness Lingua Franca’ to Facilitate Cross-Team Communication
 Engaging with external fairness teams revealed a nuanced challenge: attaining mutual understanding and converging perspectives.
-Practitioners valued the fairness teams’ feedback but sometimes struggled with its practical integration due to a few key complexities.
-Firstly, the inherent structure of RS as complex deep learning models sometimes made feedback from fairness teams difficult to implement.
-P9 observed, “sometimes…[a] challenge is fundamentally unsolvable, rooted in the deep learning models.” P1 shared, “Their feedback could be something I agree with, but I cannot solve.” Secondly, the multidisciplinary nature of fairness teams led them to use terminology that was not readily understood by technical teams, affecting inter-team communication.
-P9 commented on the influence of varied backgrounds and perspectives on understanding: “sometimes I don’t understand some phrases they used in their feedback, for example, ‘inductive bias from recommender systems,’ and need to ask for more explanations.” P6 added, “Some of my colleagues are from the field of psychology and communication, and some of those terms I didn’t know before.” Thirdly, practitioners desired clearer guidance from fairness teams, particularly regarding model development.
-P7 said, “Once they mentioned a certain aspect of my model might introduce bias, I wished to understand the reasoning.” Finally, fairness teams were often involved only after models were deployed for online testing.
-P9 suggested, “Early involvement of the fairness team would have facilitated a closer working relationship.” P2 noted, “Sometimes there’s a lag in syncing up; feedback might refer to previous versions of our recommender system designs.”
-These challenges starkly contrast with how RS practitioners interact with legal teams.
+Participants valued fairness teams’ feedback but sometimes struggled to integrate it due to a few key complexities.
+First, the complexity of RS sometimes made feedback from fairness teams difficult to implement, as highlighted by P1: “Their feedback could be something I agree with, but I cannot solve,” and by P9: “Sometimes…[a] challenge is fundamentally unsolvable, rooted in the deep learning models.” Second, the multidisciplinary nature of fairness teams led them to use terminology that was not readily understood by technical teams, affecting inter-team communication:
+P6: Some of my colleagues are from the field of psychology and communication, and some of those terms I didn’t know before.
+P9: Sometimes I don’t understand phrases they use in their feedback, for example, ‘inductive bias from recommender systems,’ and need to ask for more explanations.
+Third, participants desired clearer guidance from fairness teams, particularly regarding model development.
+For example, P7 explained that, “Once [the fairness team] mentioned a certain aspect of my model might introduce bias, I wished to understand the reasoning.” Finally, fairness teams were often involved only after models were deployed for online testing.
+As P9 put it, “I could work with [fairness] closer if they could have joined [our project] earlier.”
+We are not the first to identify communication challenges as a barrier to assessing and improving fairness in practice: notably, Deng et al.
+(2023b) interviewed AI practitioners to understand “cross-functional collaboration” in industry.
+They found that practitioners addressed communication challenges by taking on “bridging roles,” in which they championed fairness efforts by initiating cross-team meetings and developing educational resources for their colleagues.
+While beneficial to their teams, practitioners who took on bridging roles struggled with under-recognition and unreasonable expectations, as their existing full-time jobs did not typically include explicit fairness-related responsibilities (Deng et al., 2023b).
+Interestingly, even though the RS practitioners we interviewed all worked with dedicated fairness teams—theoretically reducing the need for informal bridging roles—they still struggled with cross-team communication.
+Importantly, however, cross-team communication challenges were not ubiquitous: participants overwhelmingly drew a sharp contrast between their interactions with fairness teams and their interactions with legal teams.
 Due to the sensitive nature of end-user data, practitioners regularly interact with their legal team to confirm the scope of data that can be gathered.
-With increasing surveillance from authorities and global concerns over privacy on RS, one might expect the relationship between practitioners and internal policymakers to be fraught.
-However, practitioners reported that communication with legal teams was clear and direct.
-Per P2, “If they said ‘no,’ then it is a clear ‘no,’ and we know the reasons and risks behind their decision.” Similarly, P3 stated, “Their guidelines are very clear, particularly when compared with fairness guidelines, which we don’t really have.” Clear and open guidelines from legal teams made it easier for practitioners to work with them, and most practitioners acknowledged the role of the legal team in avoiding risks.
-For instance, P2 mentioned that the “legal team serves as the gatekeeper to keep us away from legal risks, and help us comply well with our company’s vision and values.” P1 summarized the issue succinctly: “Our collaborations [with fairness teams] could be streamlined if there were guidelines as clear as those from legal teams.”
+With increasing surveillance from authorities and global concerns over privacy in RS, one might expect the relationship between practitioners and internal policymakers to be fraught.
+However, participants reported that communication with legal teams was clear and direct.
+Per P2, “If they said ‘no,’ then it is a clear ‘no,’ and we know the reasons and risks behind their decision.” Similarly, P3 stated, “Their guidelines are very clear, particularly when compared with fairness guidelines, which we don’t really have.” Clear and open guidelines from legal teams made it easier for practitioners to work with them, and most participants acknowledged the role of the legal team in avoiding risks.
+P1 summarized the issue succinctly: “Our collaborations [with fairness teams] could be streamlined if there were guidelines as clear as those from legal teams.” Another thing that the technical practitioners we interviewed appreciated about legal teams was their early engagement—legal teams become involved in fairness work during the offline prototyping phase, while fairness teams typically do not engage until the online collaborative phase.
+Friction in cross-team understanding was a hindrance:
+P9: Sometimes, [the fairness team asks] for a lot of abbreviations…which can be found in our review documents.
+Personally, I could understand that joining in the middle of a project, it is hard for them to know a lot of discussions and terms that we used.
+Similarly, P2 felt that fairness teams did not always provide actionable recommendations, saying “[Fairness is] sometimes not on the same page—they gave feedback based on outdated [RS] designs.” These observations suggest potential paths forward for improving communication between fairness and technical teams.
 7. Calls to Action
-In this section, we map the pain points uncovered in our interviews to calls for action, targeted towards RS and fairness communities alike.
-7.1. Building Institutional Knowledge Through Documentation and Collaboration
+In this section, we map the challenges uncovered in our interviews to calls for action, highlighting both where big tech companies and practitioners are responsible for implementing evidence-based changes to better integrate fairness into their RS workflows as well as where novel HCI research can support real-world fairness efforts.
+7.1. Building Institutional Knowledge Through Documentation
 A key challenge reported by RS practitioners was adapting fairness metrics that had been previously developed at their organizations for new fairness work.
-Teams are composed of individuals from a wide variety of personal backgrounds, numerous project stages are spread over time, and systems are improved dynamically after deployment, when fairness effects may be hard to predict and can potentially self-reinforce in feedback loops.
-Due to these challenges, documenting prior fairness efforts is critical for building institutional fairness knowledge.
+Teams are composed of individuals from a wide variety of personal backgrounds, numerous project stages are spread over time, and systems are modified dynamically after deployment, when fairness effects may be hard to predict and can potentially self-reinforce in feedback loops.
+Due to these challenges, documenting fairness work is critical for building institutional fairness knowledge.
 However, as multiple practitioners pointed out, while there are standard practices for documenting and approving code changes, there is comparably little infrastructural support for documenting approaches to assess and improve fairness.
-Researchers have previously called for such fairness documentation (Heger et al., 2022), e.g., in the style of datasheets for datasets (Gebru et al., 2021; Papakyriakopoulos et al., 2023).
-An internal template for fairness documentation for RS can help to improve the process in the immediate term and also build institutional knowledge.
-Considering that RS involve massive amounts of content, large user bases, rapidly changing environments, and large teams of developers, it is surprising that the practitioners we interviewed did not identify stronger support for maintaining team cohesion and situational awareness during project development.
-While collaboration was explicitly cited as a stage of the fairness process by practitioners, practitioners did not report receiving specific technical or procedural support to facilitate cross-team collaboration.
-We believe there is a need for research into best practices for collaboration and information exchange during RS development to better support cross-team fairness efforts.
-7.2. Encouraging Researcher and Organizational Investment in Multi-Stakeholder Fairness
-To tackle the challenges involved in improving fairness in RS, organizations must invest intellectual resources into researching, or working to implement academic research on, several key areas.
-Firstly, it is necessary to develop more comprehensive optimization approaches, considering multiple objectives, to balance (or avoid) tradeoffs between different stakeholder groups.
-Secondly, further investigations into group-wise fairness performance comparisons should address the “too many groups” challenge in RS by making it easier to assess fairness across a very high numbers of groups.
-Similarly, there should also be an increased focus on intra-group fairness to align with multi-faceted and intersectional conceptions of diversity in the real world.
-Additionally, researchers and organizations should work to develop standardized guidelines for making tradeoff decisions to scaffold decision-making when fairness tradeoffs between stakeholder groups is unavoidable.
-This would help avoid the ad hoc and potentially arbitrary decision-making that currently occurs when RS practitioners are forced to make these decisions in real time as tradeoffs arise.
+The Role of Practitioners and Organizations
+Practitioners and organizations must document their fairness metrics and prior fairness work.
+Practitioners and organizations are also responsible for effectively building on institutional knowledge.
+For example, the practitioners we interviewed reported that, when available, they re-used metrics from prior projects in the initial prototyping stage of the fairness workflow, but typically developed around five brand-new metrics by the end of the collaboration stage.
+This ad-hoc approach impedes the accumulation of institutional knowledge.
+It may harm procedural fairness as well, as different fairness standards are applied across different systems, and standards may not always be rigorously evaluated for validity.
+Thus, we suggest that practitioners should design metrics with extensibility, or the ability for metrics to be “adapted for different systems, use cases, and deployment contexts” in mind (Harvey et al., 2025).
+The Role of HCI Researchers
+HCI and ML fairness researchers have previously led calls for improved documentation in both research and industry.
+Researchers have proposed documentation frameworks including datasheets for datasets (Gebru et al., 2021) and model cards for model reporting (Mitchell et al., 2019)—and have led efforts to understand the extent to which documentation meets the needs of and is adopted by industry practitioners (Heger et al., 2022; Boyd, 2021).
+More recently, researchers have begun to propose domain-specific augmentations to documentation frameworks, such as datasheets for speech datasets (Papakyriakopoulos et al., 2023) or for digital cultural heritage datasets (Alkemade et al., 2023).
+Others have documented how AI practitioners themselves customize and adapt artifacts from research, like fairness checklists, to suit their needs and align with their workflows (Madaio et al., 2024; Morley et al., 2023; Madaio et al., 2020).
+Going forward, HCI researchers are uniquely well-suited to determining how such artifacts should be designed, i.e., which design practices lead to the development of adaptable or extensible documentation frameworks.
+Some researchers are already working on this: for example, Madaio et al.
+(2024) suggest that designers of artifacts intended for use by industry practitioners embrace “positive ambiguity,” or under-specification, so that artifacts can more easily be adapted.
+Similarly, Harvey et al.
+(2025) propose that artifacts should be designed to be modifiable and modular in order to achieve extensibility.
+Both Madaio et al.
+(2024) and Harvey et al.
+(2025) leave open the question of how ambiguous or modifiable documentation frameworks and other artifacts can be before they no longer perform their intended function of standardization, suggesting one clear avenue for future HCI research.
+7.2. Investing in Multi-Stakeholder Fairness
+Multi-stakeholder fairness considerations raised a variety of challenges, including managing conflicting stakeholder interests; prioritizing between many different, non-conflicting, types of fairness; and managing the technical and prioritization challenges associated with handling high numbers of user groups.
+The Role of Practitioners and Organizations
+To tackle the challenges involved in improving fairness in RS, organizations must invest resources into researching, or working to implement academic research on, several key areas.
+First, it is necessary to develop more comprehensive optimization approaches, considering multiple objectives, to balance (or avoid) tradeoffs between different stakeholder groups.
+Second, research should address the “too many groups” challenge in RS by making it easier to assess fairness across a very high numbers of groups.
+Similarly, there should be an increased focus on intra-group fairness to align with multi-faceted and intersectional conceptions of diversity in the real world (Wang, 2025; Wang et al., 2022; Kanubala and Valera, 2025).
+While ML researchers can support this research, the massive scale of modern RS mean that much of this work will be computationally and financially expensive, and thus much of the responsibility should fall to well-resourced big tech organizations who also profit from RS (Widder et al., 2024).
+The Role of HCI Researchers
+HCI researchers can advance multi-stakeholder fairness by developing transparent and standardized guidelines to scaffold decision-making when fairness tradeoffs between stakeholder groups are unavoidable.
+Value-sensitive design (VSD) methods, which embrace tensions between conflicting stakeholder values as a core theoretical construct, could prove a valuable approach (Friedman and Hendry, 2019; Boyd, 2022; Zhu et al., 2018; Simon et al., 2020; Shen et al., 2022).
+Both Zhu et al.
+(2018) and Simon et al.
+(2020) have suggested that VSD’s methodology of iterative conceptual, empirical, and technical investigations can be incorporated into the design of fair algorithms.
+As a concrete example, Shen et al.
+(2022) use a VSD approach to create the Model Card Authoring Toolkit, which is intended to help community members with “tradeoff deliberation” while designing decision-making algorithms.
+We suggest that HCI research focused on similar VSD-informed efforts could help avoid the ad hoc and potentially arbitrary decision-making that currently occurs when RS practitioners are forced to make fairness tradeoffs in real time.
 Guidelines should be developed in collaboration with diverse stakeholder groups, and should encourage practitioners to question, rather than reinforce, existing power dynamics on RS platforms.
 7.3. Integrating Fairness Considerations into Workflows Early and Often
+RS practitioners typically did not interact with fairness teams until models had been deployed.
+This delay caused friction, as fairness teams sometimes had out-of-date information on technical terminology or system updates.
+At the same time, during online phases of model development (when fairness teams were in the loop), technical teams were consistently “on-call” (P2) and reported prioritizing maintaining model uptime over fairness fixes.
+The Role of Practitioners and Organizations
 On an organizational level, we found that practitioners encountered competing incentives for integrating fairness into their workflows.
-While practitioners’ primary job responsibilities were to create and manage RS, every practitioner interviewed felt both a personal interest and an explicit requirement to assess and improve fairness in their models.
+While practitioners’ primary job responsibilities are to create and manage RS, every practitioner interviewed felt both a personal interest and an explicit requirement to assess and improve fairness in their models.
 At the same time, however, we noticed many cases where practitioners had to backtrack or iterate on deployed models in order to mitigate unexpected instances of unfairness.
-Had they invested additional time during project formation and received support from fairness teams earlier, perhaps such issues could have been detected when changes would have been easier to make.
-Our workflow models (Figures 1 and 2) highlight the opportunity for early engagement with fairness teams, at transition points in the workflow and during phases of the RS workflow that align with phases of the fairness process.
-We also call for additional research into the organizational context of RS in order to identify organizational barriers beyond the scope of our investigation.
+Had they invested additional time during project formation and received support from fairness teams earlier, perhaps such issues could have been detected when changes were easier to make.
+Our workflow model (Figures 1) highlights the opportunity for earlier engagement with fairness teams.
+The practitioners we interviewed explicitly stated that they would have preferred fairness teams to be involved with their projects earlier on; there is ample opportunity for this during the prototyping and internal phases of the fairness workflow.
+The Role of HCI Researchers
+Researchers are well-equipped to develop tools that support industry practitioners in incorporating fairness into their workflows.
+For example, Boenig-Liptsin et al.
+(2022) have proposed the Ethos Lifecycle, an interactive online tool designed to increase the extent to which data science practitioners consider positionality, power, the interrelatedness of social and technical considerations, and the narratives animating their work.
+The tool is built around a model of a data science workflow and invites users to consider relevant ethical questions at each stage of the workflow.
+Similarly, Boyd (2022) uses VSD principles to create a field guide to inform practitioners about technical mitigations to ethical issues that they identify in the course of their work.
+That research was grounded in the concept of ethical sensitivity, which Boyd and Shilton (2021) define as “the moment of noticing an ethical problem (recognition), the process of building understanding of the situation (particularization), and the decision about what to do (judgment).” Following Boyd and Shilton (2021), we suggest that additional HCI research is needed to understand the ways in which ethical recognition can be triggered and how such triggers can be built in to industry workflows.
+Furthermore, we suggest that HCI researchers can build on our model of the RS practitioner workflow to create teaching tools that could improve ethical particularization and judgment and encourage early consideration of fairness in the development of RS.
 7.4. Developing a ‘Fairness Lingua Franca’
-Across Industry Teams
-We identified cases where poor communication between fairness teams and technical practitioners led to confusion and even deprioritization of fairness efforts.
+We identified several cases where poor communication between fairness teams and technical practitioners led to confusion and even deprioritization of fairness efforts.
+The Role of Practitioners and Organizations
 While legal teams offered explicit guidance and directives to practitioners, the nature of fairness feedback—reliant on nuance and not necessarily immediately applicable to complex models—proved much harder to parse and operationalize in systems.
@@
 Practitioners reported that fairness teams came from multiple disciplines, including some (like psychology and communication) which may not share common terminology with ML.
-RS practitioners lacked the broad interdisciplinary expertise represented on fairness teams, resulting in a disconnect in fully understanding the implications and context of fairness feedback.
-The dynamic environment of RS exacerbated this gap, leading practitioners to optimize by solving easy issues while neglecting other—potentially more serious—issues.
-We found that the academic literature was often viewed as a reliable, albeit not universally applicable, resource by both practitioners and fairness teams.
-We therefore posit that HCI and ML researchers might also help with the necessary ‘translation’ work to provide useful vocabulary and common ground for practitioners.
-Likewise, embedding fairness practices more deeply in technical teams through additional roles, building domain expertise in mutual fields, and pursuing techniques for more consistent documentation and information flow (especially at the beginning of the RS workflow) might also help to establish common ground.
-Across Industry and Academia
-We found that knowledge from research communities not only shaped many practitioners’ awareness of fairness but also gave them some basic principles to follow.
-Yet practitioners also regularly developed their own ad hoc fairness metrics instead of using those described in academic literature.
+RS practitioners lacked the broad interdisciplinary expertise represented on fairness teams, resulting in barriers to fully understanding the implications and context of fairness feedback.
+The dynamic environment of RS exacerbated this gap, leading practitioners to optimize by solving easy issues while neglecting other—potentially more serious—ones.
+Nevertheless, the gap could be reduced by more frequent meetings between technical and fairness teams in which shared terminologies are established.
+Likewise, embedding fairness practices more deeply in technical teams through additional roles, building domain expertise in mutual fields, and pursuing techniques for more consistent documentation and information flow (especially at the beginning of the RS workflow) might help to establish common ground.
+We found that, when participants were aware of RS fairness research, research could provide basic principles for practitioners to follow or represent starting points for bespoke metrics.
+Yet several practitioners reported that they did not have the time to regularly review academic research, and thus that they consistently developed their own ad hoc fairness metrics instead.
 This creates a patchwork of practices that at times adheres to academic recommendations and at other times drifts far from them.
-While practitioners do not necessarily seek to disregard academic best practices, practical constraints sometimes make it hard to follow, or even known about, them.
-We call for greater collaboration between industry and academia, acknowledging that accomplishing this goal is challenging and limited by many factors beyond the scope of our study.
-However, there are clear benefits to both sides: industry practitioners may benefit from investigating novel approaches, while research communities can potentially gain exposure to new use cases and novel empirical evidence.
+While practitioners do not necessarily seek to disregard academic best practices, practical constraints sometimes make it hard to follow, or even know about, them.
+We suggest that the big tech companies that employ RS practitioners are responsible for incentivizing practitioners to devote the time necessary to stay up-to-date on academic research, i.e., by making it an explicit job responsibility as opposed to a form of invisible labor (Deng et al., 2023b).
+The Role of HCI Researchers
+We suggest that HCI researchers might help with the necessary translation work to provide common ground for technical and fairness practitioners.
+For example, Morley et al.
+(2021) propose Ethics as a Service, as a model for helping practitioners “translate principles into practice.” Ethics as a Service suggests concrete roles for organizations and practitioners to adopt organization-wide ethical principles and determine the appropriate approaches for implementing them.
+Similarly, Delgado et al.
+(2022) identify factors that led to the successful translation of real-world legal requirements into an AI system through a participatory design task, including extended co-design and evaluation of the AI solutions and the creation of opportunities for expertise to “flow in multiple directions.” HCI researchers can build on this prior work to identify best practices for collaboration and information exchange during RS development to better support cross-team fairness efforts.
+A particularly useful starting point for this work is Deng et al.
+(2023b), who describe how practitioners take on “bridging” roles to build AI fairness expertise at their organizations and identify a variety of efforts undertaken by practitioners with fairness expertise to educate practitioners without that expertise, including hosting design workshops and hackathons and developing guidebooks and toolkits.
+At the same time, the work finds that bridging efforts are labor-intensive and under-appreciated (Deng et al., 2023b).
+Thus, we call for additional research into the effectiveness of different bridging strategies, as well as how bridging work can be made easier for practitioners.
 8. Discussion
-8.1. The Path Forward: RS Practitioners Emphasize Benefits of Fairness Work Despite Challenges
 Despite the aforementioned challenges, we have reason to be optimistic: RS practitioners consistently ascribe importance to fairness work.
@@
 While most practitioners reported that balancing daily work and fairness considerations was somewhat stressful, most also believed that fairness is an important and necessary part of their work.
-P9 described incorporating fairness as “…tiring but really rewarding.” Practitioners pointed out that fairness felt particularly important given the uniquely large impact that RS play on society.
-P1 notes: “It is tiring; however, I believe this is the right thing to do.
+P9 described incorporating fairness as “tiring but really rewarding.” Practitioners pointed out that fairness felt particularly important given the uniquely large impact that RS play on society:
+P1: It is tiring; however, I believe this is the right thing to do.
 It is important to remove bias as much as possible as our system serves millions of users worldwide.
-Discriminating behaviors imposed by our system might greatly affect their life.” Practitioners expressed a desire to improve the world through their systems and identified strong intrinsic motivations for improving fairness, even if it is arduous.
-Some practitioners cited public and academic discourse on fairness as motivating; for instance, P1 mentioned, “Thorough discussions from existing papers, you know, like COMPAS, about how unfairness affects other people’s life greatly shaped my view on this.
-As I have the chance to repair this, I will do my best—though it is not easy.”
+Discriminating behaviors imposed by our system might greatly affect their lives.
+Practitioners expressed a desire to improve the world through their systems and identified strong intrinsic motivations for improving fairness, even if it is arduous.
+Some practitioners cited public and academic discourse on fairness as motivating:
+P1: Thorough discussions from existing papers, like COMPAS (Angwin et al., 2016), about how unfairness affects other people’s life greatly shaped my view on this.
+As I have the chance to repair this, I will do my best—though it is not easy.
 8.1. Limitations and Ethical Considerations
@@
 In particular, our findings may not generalize to RS practitioners working on smaller-scale RS.
-Similarly, this paper focuses on machine learning practitioners working on RS who are leading engineering or research efforts; we did not interview members of legal, data, or fairness teams.
-The desires we uncovered may also not be fully actionable: there is a fundamental tension between the desire for fairness rules that function similarly to rigid and straightforward legal and data governance requirements, and the ability actually to implement such rules in the wild (Ali et al., 2023).
-Finally, our semi-structured interviews might be limited by the remote setting, potential risks of exposing corporate secrets, or participant self-presentation maintenance.
+Similarly, it is likely that fairness is more salient to the specific subset of individuals who agreed to be interviewed than it would be to a representative sample of RS practitioners.
+This does not necessarily imply that the individuals we interviewed feel differently about fairness than the average RS practitioner, but it is certainly possible that they were more likely to believe that fairness was “rewarding” (P9) and “the right thing to do” (P1) (see §8).
+In any case, we do not attempt to draw generalizations about how common the fairness challenges and beliefs we identify are among RS practitioners as a whole—we only claim that these challenges exist, and that they impact the development of large-scale RS.
+Furthermore, this paper focuses on technical practitioners working on RS who are leading engineering or research efforts; we did not interview members of legal, data, or fairness teams.
+We chose to do this because we were specifically interested in how practitioners who are actively involved in building RS incorporate fairness into their responsibilities.
+Nevertheless, this does prevent us from analyzing the extent to which the workflows described by legal, data, and fairness teams align with—and deviate from—those described by technical teams.
+Further, our semi-structured interviews might be limited by the remote setting, potential risks of exposing corporate secrets, or participant self-presentation maintenance.
 While we found participants to generally be quite frank about their experiences, it is possible that they provided a more critical or a rosier picture of the environment due to their perceived expectations.
+Finally, the challenges we uncovered are likely not fully addressable: there is a fundamental tension between the desire for fairness rules that function similarly to rigid and straightforward legal and data governance requirements, and the ability to actually implement such rules in the wild (Ali et al., 2023).
 Ethical Considerations
@@
 Furthermore, there is a risk that the emphasis on certain strategies might overshadow other equally important aspects of fairness in RS.
-We recognize the importance of ongoing critical evaluation and discourse in the community to mitigate these potential negative impacts and encourage a diverse range of perspectives and solutions in addressing structural issues of fairness in recommender systems.
+We recognize the importance of ongoing critical evaluation and discourse in the community to mitigate these potential negative impacts and encourage a diverse range of perspectives and solutions when addressing structural issues of fairness in RS.
 9. Conclusion
-In this work, we investigated fairness practices in RS through a semi-structured interview study of 11 technical practitioners who are building recommender systems.
-We constructed a process-level map of RS practitioners’ daily workflow, focusing on how they interact with legal, data, and fairness teams at their organizations and on how they incorporate fairness into their existing practices.
+In this work, we investigated fairness practices in RS through a semi-structured interview study of 11 technical practitioners who are building RS.
+We constructed a map of RS practitioners’ daily workflow, focusing on how they interact with legal, data, and fairness teams at their organizations and on how they incorporate fairness into their existing practices.
 We identified key technical challenges to incorporating fairness, such as defining fairness in the context of RS and navigating the complexity introduced by RS; namely, multi-stakeholder fairness considerations and dynamic environments.
-We also identified organizational challenges resulting from cross-team interactions, including a lack of time to conduct fairness work and an communication challenges across teams stemming from a lack of a ‘fairness lingua franca.’ We suggest that building institutional knowledge by improving documentation and collaboration practices, encouraging researcher and organizational investment into multi-stakeholder fairness approaches, integrating fairness considerations into RS workflows early and often, and developing a ‘fairness lingua franca’ can lead to improvements in RS fairness.
+We also identified organizational challenges resulting from cross-team interactions, including a lack of time to conduct fairness work and communication challenges across teams.
+We suggest that building institutional knowledge by improving documentation practices, encouraging investment into multi-stakeholder fairness approaches, integrating fairness considerations into RS workflows early and often, and developing a ‘fairness lingua franca’ can lead to improvements in RS fairness.
 Acknowledgements.
@@
 We would like to thank the associate chairs and anonymous reviewers for their invaluable feedback.
-We also would like to thank Dataprep team from Simon Fraser University for the valuable input of mapping of common functions and user intentions.
+We also would like to thank the Dataprep team from Simon Fraser University for the valuable input of mapping common functions and user intentions.
 All opinions, findings, and conclusions in this paper are those of the authors and do not necessarily reflect the views of the funding agencies.
+Appendix A Semi-Structured Interview Guide
+Interview Phase Question Type Questions
+Ice-Breaking Participant background Please tell me a bit more about yourself.
+Participant company Please tell me something about the recommender system that you are helping to build.
+Participant experience How long have you been working in this field?
+How long you have been working at your current company?
+Are you focusing on research, production, or a mixture of both?
+Workflow and Approach to Fairness Participant Workflow (opening questions) Could you describe a recent project’s workflow?
+Are there external teams that you interact with regularly?
+Could you categorize your basic daily stages?
+How do you feel about your workflow?
+Participant Workflow (follow-up questions) Which stage do you think you spend most of your time on?
+Would you mind sharing a recent interaction with [X] team?
+How do you feel while interacting with [X] team?
+Participant Approach to Fairness How do you define bias in your daily work?
+Where do you find metrics and tools to quantify and detect bias?
+Can you share a recent experience detecting bias in the RS that you work with?
+Do you rely on a taxonomy of unfairness or harm in your work? If yes, how do you find it useful?
+What amount of time do you spend on fairness in your daily work?
+How do you feel about your experience with fairness in RS?
+Organizational Approach to Fairness Is fairness in your organization a team effort, or is responsibility for fairness distributed just to one person or a small group?
+What do you think of this assignment?
+In terms of mitigation approaches, do you prefer model-centric ones or data-centric ones? Why?
+Could you describe a recent time that improving fairness failed or was extremely difficult?
+What are the barriers you find that will prevent you from improving fairness?
+How do you like the current partition of the time? Would you prefer to spend more or less effort on fairness? Why?
+Work Experience Participant Preferences Describe fairness-related tools that you are familiar with. Which tools are you more or less likely to use? Why?
+Can you rate your professional experiences in RS in terms of emotional support, work intensity, and stress level?
+Table 3.
+Semi-Structured Interview Guide.
+As is typical of semi-structured interview studies, not all participants were asked the exact same set of questions in the same order.
+Some participants were asked additional follow-up questions as appropriate.
+A semi-structured interview guide.
+It includes: (1) ice-breaking questions: Please tell me a bit more about yourself, Please tell me something about the recommender system that you are helping to build, How long have you been working in this field?, How long you have been working at your current company?, Are you focusing on research, production, or a mixture of both?
+(2) Questions about participants’ workflows and approaches to fairness: Could you describe a recent project’s workflow?, Are there external teams that you interact with regularly?, Could you categorize your basic daily stages?, How do you feel about your workflow?, Which stage do you think you spend most of your time on?, Would you mind sharing a recent interaction with [X] team?, How do you feel while interacting with [X] team?, How do you define bias in your daily work?, Where do you find metrics and tools to quantify and detect bias?, Can you share a recent experience detecting bias in the RS that you work with?, Do you rely on a taxonomy of unfairness or harm in your work?
+If yes, how do you find it useful?, What amount of time do you spend on fairness in your daily work?, How do you feel about your experience with fairness in RS?, Is fairness in your organization a team effort, or is responsibility for fairness distributed just to one person or a small group?
+What do you think of this assignment?, In terms of mitigation approaches, do you prefer model-centric ones or data-centric ones?
+Why?, Could you describe a recent time that improving fairness failed or was extremely difficult?, What are the barriers you find that will prevent you from improving fairness?, How do you like the current partition of the time?
+Would you prefer to spend more or less effort on fairness?
+Why?
+(3) Work experience questions: Describe fairness-related tools that you are familiar with.
+Which tools are you more or less likely to use?
+Why?
+Can you rate your professional experiences in RS in terms of emotional support, work intensity, and stress level?